### v2

# Global Extraction of the $\rm{}^{12}C$ Nuclear Electromagnetic Response Functions (${\cal R}_{L}$ and ${\cal R}_{T}$) and Comparisons to Nuclear Theory and Neutrino/Electron Monte Carlo Generators

Abstract
We have performed a global extraction of the ${\rm{}^{12}C}$ longitudinal (${\cal R}_{L}$) and transverse (${\cal R}_{T}$) nuclear electromagnetic response functions from an analysis of all available electron scattering data on carbon. The response functions are extracted for energy transfer $\nu$, spanning the nuclear excitation, quasielastic (QE), resonance and inelastic continuum over a large range of the square of the four-momentum transfer, $Q^{2}.$ In addition, we perform a universal fit to all ${\rm{}^{12}C}$ electron scattering data which also provides parmeterizations of ${\cal R}_{L}$ and ${\cal R}_{T}$ over a larger kinematic range. Given the nuclear physics common to both electron and neutrino scattering from nuclei, extracted response functions from electron scattering spanning a large range of $Q^{2}$ and $\nu$ also provide a powerful tool for validation and tuning of neutrino Monte Carlo (MC) generators. In this paper we focus on the nuclear excitation, single nucleon (QE-1p1h) and two nucleon (2p2h) final state regions and compare the measurements to theoretical predictions including “Energy Dependent-Relativistic Mean Field” (ED-RMF), “Green’s Function Monte Carlo” (GFMC), “Short Time Approximation Quantum Monte Carlo” (STA-QMC), an improved superscaling model (SuSAv2), “Correlated Fermi Gas” (CFG), as well as the NuWro, and achilles  generators. Combining the ED-RMF-QE-1p1h predictions with the SuSAv2-MEC-2p2h predictions provides a good description of ${\cal R}_{L}$ and ${\cal R}_{T}$ for both single nucleon (from QE and nuclear excitations) and two nucleon final states over the entire kinematic range.
pacs:
I Introduction
Electron scattering cross sections on nuclear targets are completely described by longitudinal (${\cal R}_{L}$) and transverse (${\cal R}_{T}$) nuclear electromagnetic response functions. Here ${\cal R}_{L}$ and ${\cal R}_{T}$ are functions of the energy transfer $\nu$ (or excitation energy $E_{x}$) and the square of the 4-momentum transfer $Q^{2}$ (or alternatively the 3-momentum transfer $\bf q$). Recent theoretical  Mihaila and Heisenberg (2000); Lovato et al. (2016); Cloët et al. (2016); Franco-Munoz et al. (2022, 2023); Sobczyk et al. (2020); Carlson et al. (2015, 2002) of ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ can be tested by comparing the predictions to experimental data. We perform extraction of both response functions as functions of (${\bf q},\nu$), as well as ($Q^{2},\nu)$ from an analysis of all available inclusive cross section data for $\rm{}^{12}C$, and compare them to theoretical predictions.
Given the nuclear physics common to both electron and neutrino scattering from nuclei, extracted response functions from electron scattering spanning a large range of $Q^{2}$ and $\nu$ also provide a powerful tool for validation and tuning of neutrino Monte Carlo (MC) generatorsGallagher and Hayato (2022); Workman et al. (2022). Early MC generators include neugen Gallagher (2002) and nuance Casper (2002). More recent calculations and MC generators include neut Hayato (2002); McKean (2024); Dolan et al. (2023), gibuu (Giessen Boltzmann-Uehling-Uhlenbeck) Leitner et al. (2009); Buss et al. (2012), NuWro NuW (2024), achilles  (A CHIcago Land Lepton Event Simulator) Isaacson et al. (2023) and genie Andreopoulos et al. (2010). With the advent of DUNEAbi et al. (2020)(Deep Underground Neutrino Experiment), the next generation neutrino oscillation experiments aim to search for CP violations in neutrino oscillations. Therefore, current neutrino MC generators need to be validated and tuned over the complete range of relevance in $Q^{2}$ and $\nu$ to provide a better description of the cross sections for electron and neutrino interactions.
In this paper we focus on the nuclear excitation, quasielastic (QE) and two nucleon final state (2p2h) regions and compare the extracted $\rm{}^{12}C$ ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ to nuclear theory predictions of the “Green’s Function Monte Carlo” (GFMC) Lovato et al. (2016), the “Energy Dependent-Relativistic Mean Field” (ED-RMF) Franco-Munoz et al. (2022, 2023), the “Short Time Approximation Quantum Monte Carlo” (STA-QMC) Andreoli et al. (2024), the “Correlated Fermi Gas” (CFG) Bhattacharya et al. (2024), the NuWro theoretical approach NuW (2024), the achilles Isaacson et al. (2023) theoretical approach, and an improved superscaling model (SuSAv2)Gonzaléz-Jiménez et al. (2014); Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023). We have not done comparisons to the gibuu theoretical approach because nuclear response function predictions by gibuu are not readily available.
The theoretical approaches of NuWro and achilles  have been implemented in the corresponding neutrino MC generators. Recently, ED-RMF has been implementedMcKean (2024) in an update of the neut neutrino MC generator. An earlier version of SuSAv2 (QE+2p2h) has been implemented in genie, SuSAv2-inelastic is being implemented in genie, and STA-QMC for $\rm{}^{4}He$ only has been implemented in genie. In this paper we focus on testing theoretical predictions. The testing of the implementation of the theoretical approaches into the genie neutrino MC generator (run in electron scattering mode) will be presented in a future communication. Similar investigation of electron scattering data on $\rm{}^{40}Ca$ and $\rm{}^{56}Fe$ (as well as hydrogen and deuterium data) are currently under way.
Figure 1: Nuclear excitation region: Comparison of the ${\rm{}^{12}C}$ longitudinal ${\cal R}_{L}/Z^{2}$ (left) and transverse ${\cal R}_{T}/Z^{2}$ (right) response functions extracted by Yamaguchi:71 Yamaguchi et al. (1971) (black squares) versus excitation energy $E_{x}$, to the response functions extracted from the Christy-Bodek universal fit to all available electron scattering cross section data on ${\rm{}^{12}C}$ (solid red line). The contributions from nuclear excitations with $E_{x}<12$ MeV are multiplied by (1/6). The QE contribution to the total response functions is represented by the red dashed line. The response functions for all states in the region of the Giant Dipole Resonance (20-30 MeV) region are modeled as one average broad excitation. Note there are no transverse nuclear excitations with $E_{x}<10$ MeV. (Figure adapted from  Bodek and Christy (2023)).
Figure 2: Nuclear excitation region: Comparison of our fits to published radiatively corrected inelastic electron scattering cross sections on ${\rm{}^{12}C}$ for excitation energies less than 50 MeV. The cross sections for excitation energies less than 12 MeV are multiplied by (1/6). The pink solid line is the predicted total cross section from the Christy-Bodek universal fit Bodek and Christy (2022, 2023) to all electron scattering data on ${\rm{}^{12}C}$. The fit includes nuclear excitations, a superscaling QE model Maieron et al. (2002); Amaro et al. (2005, 2020); Megias et al. (2014) with Rosenfelder Pauli suppression Rosenfelder (1980) (dashed blue line), “Transverse Enhancement/Meson Exchange Currents” (dot-dashed line) and pion production processes (at higher excitation energies). The data are from Yamaguchi:71 Yamaguchi et al. (1971) except for the cross sections for $E_{0}=54$ MeV at 180 (from Goldemberg:64 Goldemberg and Barber (1964)) and the cross sections for $E_{0}=65$ MeV at 180 (from Deforest:65 Deforest et al. (1965)). The measurements at 180 are only sensitive to the transverse form factors. (Figure adapted from  Bodek and Christy (2023)).
Figure 3: Top: Comparison of the Christy-Bodek fit to measurements. (a) A recent (not included in the fit) CLAS-e4nu:2023 cross sectionIsaacson et al. (2023) measurement at E0=1.159 GeV and $\theta$=37.450 (multiplied by 0.85). (b) A Sealock:1989 measurement Sealock et al. (1989) at 1.108 GeV and $\theta$=37.50 (multiplied by 1.06) (c) A recent (not included in the fit) Mainz:2024Mihovilovič et al. (2024) measurement at 0.855 GeV and $\theta$=700 (multiplied by 1.03). (d-f) Zeller:1973 measurementsZeller (1973a) (not included in the fit) for 2.0, 2.5 and 2.7 GeV at $\theta$=37.50 (found to be inconsistent with world data). Bottom Left: Comparison of the high resolution measurement of ${\cal R}_{T}$ in Yamaguchi:71 Yamaguchi et al. (1971) for $Q^{2}=0.085$ GeV2 (versus $E_{x}$) to values of ${\cal R}_{T}$ extracted from the lower resolution 180 electron scattering cross-section measurements published in Ryan:84 Ryan et al. (1984) (multiplied by 1.05). Bottom Right: $W^{2}$ range for the fixed $Q^{2}$ values investigated in this analysis. At low $Q^{2}$ (or $\bf q$) the contribution from nuclear excitation is significant.
Previous extractions of ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ for $\rm{}^{12}C$ were performed with a limited set of cross section data. Consequently the extractions are only available for a very limited range of $\bf q$ and $\nu$. The first extraction of ${\cal R}_{T}({\bf q},\nu)$ and ${\cal R}_{L}({\bf q},\nu)$ for $\rm{}^{12}C$ was performed in 1983 and used experimental cross sections data from only one experiment at Saclay (Barreau:83Barreau et al. (1981, 1983a, 1983b)). The extractions were performed for $\bf q$ values of 0.30, 0.40 and 0.55 GeV and values of $\nu$ in the quasielastic (QE) region. A later extraction using the same Saclay data in combination with cross sections measured at the Stanford Linear Accelerator Center (SLAC) was published in 1996 (Jourdan:96Jourdan (1996, 1995)) also for three similar values of $\bf q$ (0.30, 0.38 and 0.57 GeV) and values of $\nu$ in the QE region. The most recent extraction was published in 2021 (Buki:21Buki and Timchenko (2021)) used a limited set of cross section data measured at Kharkov. That extraction was only performed for ${\bf q}=0.3$ GeV in the QE region and has larger uncertainties than the two earlier analyses. An extraction of ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ for $\rm{}^{12}C$ at $Q^{2}$=0.1 GeV2 in the $\Delta$(1232) nucleon resonance region was published in 1998 by Baran and collaborators Baran et al. (1988).
In this communication we report on the extraction of ${\cal R}_{L}$ and ${\cal R}_{T}$ for ${\rm{}^{12}C}$ at 18 distinct values of both q and $Q^{2}$ by including all available electron scattering cross section measurements on carbon. We note that a significant number of cross section measurements were not available in tabular form and required digitization of the cross sections from figures in the publications. As summarized in Table 1, the response functions are extracted for a large range of energy transfer $\nu$ at 18 fixed values of $\bf q$ in the range $3.750.1<{\bf q}<$ 3.75  GeV and at 18 fixed values of $Q^{2}$ in the range $3.50<Q^{2}<$ 3.5 GeV2. The range of $\nu$ spans the nuclear excitation, QE, resonance and inelastic continuum. The complete data set consists of about 10,000 ${\rm{}^{12}C}$ differential cross section measurements points as well as photo-absorption data ($Q^{2}=0$). We also include high precision cross section measurements from Jefferson Lab Hall C experiment E-04-001 Alsalmi et al. (2024); Albayrak et al. (2024); Alsalmi (2019).
Relative normalization factors for each experiment are determined from a global fit to the cross section data. This is done by including the normalizations as free parameters in the fit (including the normalization uncertainty quoted for each experiment). Leveraging both the expanded data set and the global fit for centering the data to fixed q ($Q^{2}$) allowed for more precise extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ for ${\rm{}^{12}C}$. In the analysis, the global fit used all the cross section points measurements, and the individual ${\cal R}_{L}$ and ${\cal R}_{T}$ extractions used 8,500 cross section points measurements. This is because individual ${\cal R}_{L}$ and ${\cal R}_{T}$ extractions require data at both small angles and large angles for the same $\bf q$ and $\nu$ bin.
In the nuclear excitation region (excitation energy less than 50 MeV) we extract ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from fits to ${\rm{}^{12}C}$ nuclear excitation form factors by Bodek and Christy Bodek and Christy (2023). In addition, for $4016<E_{x}<$ 40 MeV, values of ${\cal R}_{L}({\bf q},E_{x})$ and ${\cal R}_{T}({\bf q},E_{x})$ are available from the analysis of Yamaguchi:71 Yamaguchi et al. (1971) for five values of $\bf q$ (0.148, 0.167, 0.205, 0.240 and 0.307 GeV). Comparisons of the fit to the Yamaguchi response functions and cross sections are shown in Figures 1 and 2, respectively.
I.1 Overview of this analysis
The Christy-Bodek universal fit to all available cross section measurements for ${\rm{}^{12}C}$ (and Hydrogen and Deuterium) is described in two recent publications Bodek and Christy (2022, 2023) (similar fits were also done for other nuclei). The fit includes nuclear elastic form factors, nuclear-excitations, quasielastic scattering, resonance production and inelastic continuum. In the quasielastic region the fit is based on the $\psi^{\prime}$ SuSA (Super Scaling Approximation) formalism Maieron et al. (2002); Amaro et al. (2005, 2020); Megias et al. (2014) with Rosenfelder Rosenfelder (1980) Pauli suppression. For the QE region the fit also includes parametrizations of a $multiplicative$ “Longitudinal Quenching Factor” at low $\bf q$ and an $additive$ “Transverse Enhancement” contribution, which is significant at intermediate $\bf q$. In the inelastic resonance and continuum region the fit is based on a Gaussian Fermi motion smeared free proton and neutron cross sections with a multiplicative medium modification factor. The relative normalizations of different ${\rm{}^{12}C}$ cross section data sets are also extracted from the fit. In summary the fit includes:
  1. 1.
    All available electron scattering data on $\rm{}^{1}H$, $\rm{}^{2}H$, $\rm{}^{12}C$ (for $\rm{}^{12}C$ we digitized additional data to supplement the data included in the QE Day (2004) and resonance Res archives as summarized in Table 2).
  2. 2.
    Coulomb corrections Gueye et al. (1999) using the Effective Momentum Approximation (EMA) in modeling scattering from nuclear targets. We use $V_{eff}$= 3.10$\pm$ 0.25 MeV for $\rm{}^{12}C$
  3. 3.
    Updated $\rm{}^{12}C$ nuclear elastic form factorBodek and Christy (2023).
  4. 4.
    Parameterizations of $\rm{}^{12}C$ nuclear excitations form factorsBodek and Christy (2023).
  5. 5.
    Superscaling function $FN(\psi^{\prime})$ parameters are re-extracted including the Fermi broadening parameter $K_{F}$.
  6. 6.
    Parameterizations of the free nucleon form factors Bosted (1995) are re-derived from all $\rm{}^{1}H$ and $\rm{}^{2}H$ data.
  7. 7.
    Rosenfelder Pauli suppression Megias et al. (2014); Rosenfelder (1980) which reduces and modifies the QE distribution at low $\bf{q}$ and $\nu$.
  8. 8.
    Updates of fits Bosted and Mamyan (2012) to inelastic electron scattering data (in the nucleon resonance region and inelastic continuum) for $\rm{}^{1}H$ and $\rm{}^{2}H$, providing the structure functions for the proton and neutron.
  9. 9.
    A $\bf q$ dependent $E_{\mathrm{shift}}^{QE}(\bf q)$ parameter for the QE process to account for the optical potential Bodek and Cai (2019, 2020) of final state nucleons. The separation energies of a proton and a neutron from $\rm{}^{12}C$ are 16 MeV and 18.7 MeV, respectively. Therefore, we ensure that the QE cross section on protons is zero for excitation energy below 16 MeV, and the QE cross section on neutrons is zero for excitation energy less than 18.7 MeV.
  10. 10.
    Photo-production data in the nuclear excitation region, nucleon resonance and inelastic continuum Bosted and Christy (2008).
  11. 11.
    Gaussian Fermi motion smeared nucleon resonance and inelastic continuum Christy and Bosted (2010); Bosted and Christy (2008). The $K_{F}$ parameters for pion production and QE can be different.
  12. 12.
    Parametrizations of the medium modifications of both the inelastic ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$  structure functions responsible for the EMC effect (nuclear dependence of inelastic structure functions). These are applied as multiplicative factors to the free nucleon cross sections prior to application of the Fermi smearing.
  13. 13.
    For QE scattering we include parameterizations of additive Transverse Enhancement TE(${\bf q},\nu$) and multiplicative Longitudinal Quenching factor $F_{quench}^{L}({\bf q})$.
  14. 14.
    We include QE data at all values of $Q^{2}$ down to $Q^{2}=0.01$ GeV2 ($\bf q$=0.1 GeV) (which were not included in the Bosted-Mamyan fit Bosted and Mamyan (2012).)
  15. 15.
    The relative normalizations between different experiments are extracted from the fit.
  16. 16.
    Data sets which are inconsistent Zeller (1973b); Heimlich et al. (1974) with the world’s data are identified and not included in the analysis (see Table 2).
The primary purpose of the fit is to model cross sections used in calculation of radiative corrections in electron scattering experiments. The fit describes all electron scattering data on $\rm{}^{12}C$ and the fit’s ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are valid for a larger kinematic range than the individual ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ measurements. Therefore, the fit can also be used to validate nuclear models and tune MC generators for electron and neutrino scattering experiments over a large kinematic range.
The use of the Christy-Bodek universal fit Bodek and Christy (2022, 2023) in identifying the few data sets that are inconsistent with all the other measurements is illustrated in the top six sub-figures of Fig. 3. Here, the comparison of the fit to recent CLAS-e4nu:2023 cross sectionIsaacson et al. (2023) measurement (which was not included in the universal fit) at E0=1.159 GeV and $\theta$=37.50 shown in sub-figure (a) indicates that normalization of this data relative to all previous data is 0.85. A comparison to the Sealock:1989 measurement Sealock et al. (1989) at 1.108 GeV and $\theta$=37.50 (which was included in the universal fit) shown in sub-figure (b) indicates a normalization of 1.06. A comparison to a recent Mainz:2024Mihovilovič et al. (2024) measurement at 0.855 GeV and $\theta$=700 (which was not included in the universal fit) shown in sub-figure (c) indicates a normalization of 1.03. Comparisons to Zeller:1973 measurementsZeller (1973a); Heimlich et al. (1974) for incident energies of 2.0, 2.5 and 2.7 GeV at $\theta$=37.50 (without any normalizations) are shown in sub-figures (d-f). The Zeller:1973 measurements are excluded from the universal fit because of inconsistent normalizations between the three energies, and an unexplained shift in $\nu$ in the 2.7 GeV data.
In this analysis the fit is primarily used to calculate “bin-centering” corrections as described below. Rosenbluth Rosenbluth (1950) extractions of ${\cal R}_{T}(Q^{2},\nu)$ and ${\cal R}_{L}(Q^{2},\nu)$ require cross section measurements at different angles for the same values of ${\bf q}$ and $\nu$. We bin the cross sections in fixed bins of ${\bf q}$ and use the fit for determining “bin centering” corrections to account for the small differences in ${\bf q}$ and $\nu$ of the binned cross sections measurements from different experiments. In addition to bins in fixed values of ${\bf q}$ we also perform the same analysis in bins of fixed values of $Q^{2}$. Sample Rosenbluth plots before the application of “bin centering” corrections are shown in Fig. 4.
In order to minimize the “bin centering” corrections it is critical to bin in a kinematic variable for which features (i.e. peaks) in the cross section remain fixed independent of the angle. For this reason we perform the analysis in bins of both the excitation energy and bins of the square of the invariant final state mass ($W^{2}=M^{2}+2M\nu-Q^{2}$). We use the results of the analysis in bins of excitation energy for $E_{x}$ less than 50 MeV (a region dominated by nuclear excitations) and the results of the analysis in bins of $W^{2}$ for $E_{x}$ greater than 50 MeV (a region dominated by QE scattering and pion production). Afterwards, we convert the $E_{x}$ and $W^{2}$ values in the center of each bin to the corresponding values of $\nu$ in the center of the bin.
The longitudinal and transverse response functions (defined in section II.3) in the nuclear excitation region are well described by the universal fit Bodek and Christy (2023) to measured transverse and longitudinal form factors for each nuclear excitation. The ${\rm{}^{12}C}$ ${\cal R}_{L}/Z^{2}$ and ${\cal R}_{T}/Z^{2}$ measurements published in Yamaguchi:71 Yamaguchi et al. (1971) in the nuclear excitation region above the proton separation energy (16$<E_{x}<40$ MeV) are compared to the ${\cal R}_{L}/Z^{2}$ and ${\cal R}_{T}/Z^{2}$ predictions from the fit in Figure 1.
Rosenbluth extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ using data spanning a range of angles from different experiments are valid in the QE and pion production regions. However, in the nuclear excitation region it is not valid to combine experiments with different experimental resolutions, as this will lead to structural artifacts in the extracted response functions. For example, in the bottom left panel of Fig. 3 we show a comparison of high resolution extraction of ${\cal R}_{T}$ in Yamaguchi:71 Yamaguchi et al. (1971) for $Q^{2}$=0.085 GeV2 to ${\cal R}_{T}$ extracted from the lower resolution 180 cross section data published in Ryan:84 Ryan et al. (1984).
In contrast, at higher values of excitation energy (for $E_{x}>30$ MeV) the cross sections are relatively smooth on the scale of the experimental resolutions and a Rosenbluth analysis using cross sections from different experiments can be performed.
As shown in the bottom right panel of Fig. 3 the contribution from nuclear excitation is significant at low $Q^{2}$ (or $\bf q$). For $E_{x}<$16 MeV we use ${\cal R}_{L}$ and ${\cal R}_{T}$ from overall fits Bodek and Christy (2023) to the nuclear excitation form factors (shown as the red solid line in Figure 1). For $4016<E_{x}<$ 40 MeV (when available) we use the precise (±$\pm 3\%$) Yamaguchi:71 measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$. In addition, for ${\cal R}_{T}$, we also use electron scattering data at 180 Goldemberg and Barber (1964); Deforest et al. (1965) as shown in the right top panel of Fig. 5. For $E_{x}>30$ MeV, we extract ${\cal R}_{L}$ and ${\cal R}_{T}$ from our analysis of all available electron scattering data as described below.
Center $Q^{2}$ $Q^{2}$ Center ${\cal R}_{T}(\nu={\bf q})$ $\bf q$ $\bf q$
$Q^{2}$ low high q From $\gamma^{12}C$ low high
0($\gamma^{12}C$) 0 0
0.010 0.004 0.015 0.100 0.0016$\pm$0.0004 0.063 0.124
0.020 0.015 0.025 0.148 0.0021$\pm$0.0007 0.124 0.158
0.026 0.025 0.035 0.167 0.0028$\pm$0.0006 0.158 0.186
0.040 0.035 0.045 0.205 0.0071$\pm$0.0007 0.186 0.223
0.056 0.045 0.070 0.240 0.0134$\pm$0.0009 0.223 0.270
0.093 0.070 0.100 0.300 0.0270$\pm$0.0006 0.270 0.340
0.120 0.100 0.145 0.380 0.0324$\pm$0.0005 0.340 0.428
0.160 0.145 0.209 0.475 0.0276$\pm$0.0005 0.428 0.523
0.265 0.206 0.323 0.570 0.0262$\pm$0.0008 0.523 0.609
0.380 0.322 0.438 0.649 0.0290$\pm$0.0006 0.609 0.702
0.500 0.438 0.650 0.756 0.0299$\pm$0.0010 0.702 0.878
0.800 0.650 1.050 0.991 0.0371$\pm$0.0003 0.878 1.302
1.250 1.050 1.500 1.619 0.0414$\pm$0.0010 1.302 1.770
1.750 1.500 2.000 1.921 0.0479$\pm$0.0010 1.770 2.067
2.250 2.000 2.500 2.213 0.0542$\pm$0.0020 2.067 2.357
2.750 2.500 3.000 2.500 0.0603$\pm$0.0030 2.357 2.642
3.250 3.000 3.500 2.783 0.0664$\pm$0.0030 2.642 2.923
3.750 3.500 4.000 3.500 0.0817$\pm$0.0030 2.923 4.500
Table 1: A summary of the 18 bins in $Q^{2}$ (in GeV2) and the 18 bins in $\bf q$ (in GeV). The ${\cal R}_{T}$ (in MeV-1) values for $\nu=\bf q$ are extracted from photo-absorption cross sections.
Figure 4: Sample ${\cal R}_{L}$ ${\cal R}_{T}$ Rosenbluth plots (before bin centering corrections).
Figure 5: Nuclear excitation region: Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region ($4016<E_{x}<$ 40 MeV) for ${\rm{}^{12}C}$. The measurements are from Yamaguchi:71 Yamaguchi et al. (1971) except for values of ${\cal R}_{T}(Q^{2}=0)$ which are extracted from photo-absorption data, and values of ${\cal R}_{T}(Q^{2}=0.01)$ which are extracted from electron scattering cross section measurements at 180 (published by Goldemberg:64 Goldemberg and Barber (1964) and Deforest:65 Deforest et al. (1965)).
Data Set Normalization Error
1 Barreau:83 Barreau et al. (1981, 1983a, 1983b) 0.9919 0.0024
2 O’Connell:87 O’Connell et al. (1987) 0.9787 0.0086
3 Sealock:89 Sealock et al. (1989) 1.06 0.1000
Baran:88 Baran et al. (1988) 0.9924 0.0046
5 Bagdasaryan:88 Bagdasaryan et al. (1988) 0.9878 0.0083
6 Dai:19 Murphy et al. (2019) 1.0108 0.0053
7 Arrington9:96 Arrington et al. (1996) 0.9743 0.0133
8 Day93 Day et al. (1993) 1.0071 0.0033
9 Arrington9:99 Arrington et al. (1999) 0.9888 0.0034
10 Gaskell:21 Arrington et al. (2021); Seely et al. (2009) 0.9934 0.0051
11 Whitney:74 Whitney et al. (1974); Moniz et al. (1971) 1.0149 0.0153
12 E04-001-2005:24 Alsalmi et al. (2024); Albayrak et al. (2024); Alsalmi (2019); Alsalmi et al. (2025) 0.9981 0.0067
13 E04-001-2007:24 Alsalmi et al. (2024); Albayrak et al. (2024); Alsalmi (2019); Alsalmi et al. (2025) 1.0029 0.0070
14 Gomez:74 Gomez et al. (1994); Res 1.0125 0.0149
15 Fomin:10 Fomin et al. (2010); Liang et al. (2022) 1.0046 0.0031
16 Yamaguchi:71 Yamaguchi et al. (1971) 1.0019 0.0029
17 Ryan:84 Ryan et al. (1984) (180) 1.10 0.0130
18 Czyż:63 Czyz (1963) (not used) 1.000 0.2000
19 Bounin:63 Bounin and Bishop (1963); Lovseth (1968) 1.150 0.2300
21 Antony-Spies70 Antony-Spies et al. (1970); Yamaguchi et al. (1971) (not used) 0.95 0.25
22 Goldemberg64 Goldemberg and Barber (1964)(180) 1.100 0.1000
23 Defores:t65 Deforest et al. (1965)(180) 0.9 0.1000
24 Mainz:2024Mihovilovič et al. (2024) (not used) 1.03 0.02
25 CLAS-e4nu:2023Isaacson et al. (2023) (not used) 0.85 0.02
30 Donnelly:68 Donnelly et al. (1968); Donnelly (1970) (not used) - -
31 Zeller:73 Zeller (1973b); Heimlich et al. (1974) (not used) - -
Table 2: A summary table of the $\rm{}^{12}C$ data sets used in the Christy-Bodek universal fit and in this analysis. Shown are the normalization factors and uncertainties. Data sets 18–23 ( have larger normalization uncertainty) are early data at very low ${\bf q}$. Data sets which are inconsistent with other data sets are identified as not used.
II Inclusive Electron-Nucleon Scattering
In terms of the incident electron energy, $E_{0}$, the scattered electron energy $E^{\prime}$, and the scattering angle $\theta$, the absolute value of the exchanged 4-momentum squared in electron-nucleon scattering is given by
$Q^{2}=-q^{2}=4E_{0}E^{\prime}{\sin}^{2}\frac{\theta}{2}.$ (1)
The mass of the undetected hadronic system (nucleon and pions) is
$W^{2}=M^{2}+2M\nu-Q^{2},$ (2)
and the square of the magnitude of 3-momentum transfer vector ${\bf q}$ is
${\bf q}^{2}=Q^{2}+\nu^{2}.$ (3)
Here $M$ is the mass of the proton and $\nu=E_{0}-E^{\prime}$. In these expressions we have neglected the electron mass which is negligible for the kinematic regions investigated in this paper.
For scattering from a nuclear target such as carbon, the excitation energy $E_{x}$ is given by $E_{x}=\nu-\nu_{elastic}$ where
$\nu_{elastic}=E_{0}-\frac{E_{0}}{1+2E_{0}{\sin}^{2}\frac{\theta}{2}/M_{A}},$ (4)
and $M_{A}$ is the mass of the nuclear target. For carbon $M_{A}=12u=11.178$ GeV (1 u = 931.502 MeV). Or equivalently
$E_{x}=\nu-\frac{Q^{2}_{elastic}}{2M_{A}},$ (5)
where $Q^{2}_{elastic}$ is the $Q^{2}$ for elastic scattering from a carbon nucleus for incident energy $E_{0}$ and scattering angle $\theta$.
II.1 Description in terms of longitudinal and transverse virtual photon cross sections
This description is often used in the nucleon resonance region. In the one-photon-exchange approximation, the spin-averaged cross section for inclusive electron-nucleon (or electron-nucleus) scattering can be expressed in terms of the photon helicity coupling as
$\frac{d\sigma}{d\Omega dE^{{}^{\prime}}}=\Gamma\left[\sigma_{T}(W^{2},Q^{2})+\epsilon\sigma_{L}(W^{2},Q^{2})\right],$ (6)
where $\sigma_{T}$ ($\sigma_{L}$) is the cross section for photo-absorption of purely transverse (longitudinal) polarized photons,
$\Gamma=\frac{\alpha E^{{}^{\prime}}(W^{2}-M_{N}^{2})}{(2\pi)^{2}Q^{2}ME_{0}(1-\epsilon)}$ (7)
is the flux of virtual photons, $\alpha=1/137$ is the fine structure constant, and
$\epsilon=\left[1+2\left(1+\frac{\nu^{2}}{Q^{2}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$ (8)
is the relative flux of longitudinal virtual photons (sometimes referred to as the virtual photon polarization). Since $\Gamma$ and $\epsilon$ are purely kinematic factors, it is convenient to define the reduced cross section
$\sigma_{r}={1\over\Gamma}\frac{d\sigma}{d\Omega dE^{{}^{\prime}}}=\sigma_{T}(W^{2},Q^{2})+\epsilon\sigma_{L}(W^{2},Q^{2}).$ (9)
All the hadronic structure information is, therefore, contained in $\sigma_{T}$ and $\sigma_{L}$, which are only dependent on $W^{2}$ and $Q^{2}$. In the $Q^{2}=0$ limit $\sigma_{T}(\nu,Q^{2})$ should be equal to the measured photo-absorption cross section Bezić et al. (1969); Ahrens et al. (1975); Carrasco and Oset (1992); Bianchi et al. (1996); Muccifora et al. (1999) $\sigma_{\gamma}(\nu)$ for real photons (here $\nu$ is the energy of the photon).
II.2 Description in terms of structure functions
This description is primarily used in the inelastic continuum region. In the one-photon-exchange approximation, the spin-averaged cross section for inclusive electron-nucleon (or electron nucleus) scattering can be expressed in terms of two structure functions as follows
$\frac{d\sigma}{d\Omega dE^{{}^{\prime}}}$ $=$ $\sigma_{M}[{\cal W}_{2}(W^{2},Q^{2})+2\tan^{2}(\theta/2){\cal W}_{1}(W^{2},Q^{2})]$
$\sigma_{M}$ $=$ $\frac{\alpha^{2}\cos^{2}(\theta/2)}{[2E_{0}\sin^{2}(\theta/2)]^{2}}=\frac{4\alpha^{2}E^{\prime 2}}{Q^{4}}\cos^{2}(\theta/2)$ (10)
where $\sigma_{M}$ is the Mott cross section. The ${\cal F}_{1}$ and ${\cal F}_{2}$ structure functions are related to ${\cal W}_{1}$ and ${\cal W}_{2}$ by ${\cal F}_{1}=M{\cal W}_{1}$ and ${\cal F}_{2}=\nu{\cal W}_{2}$. The structure functions are typically expressed as functions of $Q^{2}$ and $W^{2}$, or alternatively $\nu$ and $x=Q^{2}/(2M\nu)$. At large $Q^{2}$ and $\nu$ (deep inelastic region) $x$ is the fractional momentum of the nucleon carried by the struck quark. At low $Q^{2}$ there are target mass corrections. The target mass scaling variable  Nachtmann (1973); Georgi and Politzer (1976); Barbieri et al. (1976) is $\xi_{TM}$ where,
$\xi_{TM}$ $=$ $\frac{Q^{2}}{M\nu[1+\sqrt{1+Q^{2}/\nu^{2}}]}.$ (11)
The quantity $R_{\sigma LT}(x,Q^{2})$ is defined as the ratio of the longitudinal to transverse virtual photo-absorption cross sections, $\sigma_{L}/\sigma_{T}$, and is related to the structure functions by
$R_{\sigma LT}(x,Q^{2})=\frac{\sigma_{L}}{\sigma_{T}}=\frac{{\cal F}_{L}}{2x{\cal F}_{1}},$ (12)
where ${\cal F}_{L}$ is called the longitudinal structure function.
The structure functions are expressed in terms of $\sigma_{L}$ and $\sigma_{T}$ as follows:
$K$ $=$ $\frac{Q^{2}(1-x)}{2Mx}=\frac{2M\nu-Q^{2}}{2M},$ (13)
${\cal F}_{1}$ $=$ $\frac{MK}{4\pi^{2}\alpha}\sigma_{T},$ (14)
${\cal F}_{2}$ $=$ $\frac{\nu K(\sigma_{L}+\sigma_{T})}{4\pi^{2}\alpha(1+\frac{Q^{2}}{4M^{2}x^{2}})},$ (15)
${\cal F}_{L}(x,Q^{2})$ $=$ ${\cal F}_{2}\left(1+\frac{4M^{2}x^{2}}{Q^{2}}\right)-2x{\cal F}_{1},$ (16)
or
$2x{\cal F}_{1}={\cal F}_{2}\left(1+\frac{4M^{2}x^{2}}{Q^{2}}\right)-{\cal F}_{L}(x,Q^{2}).$ (17)
In addition, $2x{\cal F}_{1}$ is given by
$2x{\cal F}_{1}(x,Q^{2})$ $=$ ${\cal F}_{2}(x,Q^{2})\frac{1+4M^{2}x^{2}/Q^{2}}{1+R_{\sigma LT}(x,Q^{2})},$ (18)
or equivalently
${\cal W}_{1}(x,Q^{2})$ $=$ ${\cal W}_{2}(x,Q^{2})\times\frac{1+\nu^{2}/Q^{2}}{1+R_{\sigma LT}(x,Q^{2})}.$ (19)
II.3 Description in terms of response functions
This description is primarily used in the nuclear excitation and QE regions. The electron scattering differential cross section is written in terms of longitudinal (${\cal R}_{L}({Q^{2}},\nu)$) and transverse (${\cal R}_{T}({Q^{2}},\nu)$) nuclear response functions Caballero et al. (2010) as
$\frac{d\sigma}{d\nu d\Omega}$ $=$ $\sigma_{M}[A{\cal R}_{L}({Q^{2}},\nu)+B{\cal R}_{T}({Q^{2}},\nu)],$ (20)
where $\sigma_{M}$ is the Mott cross section (10), $A=(Q^{2}/{\bf q}^{2})^{2}$ and $B=\tan^{2}(\theta/2)+Q^{2}/2{\bf q}^{2}$.
The relationships between the nuclear response functions, structure functions and virtual photon absorption cross sections are:
${\cal R}_{T}({\bf q},\nu)$ $=$ $\frac{2{\cal F}_{1}({\bf q},\nu)}{M}=\frac{K}{2\pi^{2}\alpha}\sigma_{T},$ (21)
${\cal R}_{L}({\bf q},\nu)$ $=$ $\frac{{\bf q}^{2}}{Q^{2}}\frac{{\cal F}_{L}({\bf q},\nu)}{2Mx}=\frac{{\bf q}^{2}}{Q^{2}}\frac{K}{4\pi^{2}\alpha}\sigma_{L},$ (22)
where ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ have units of $M^{-1}$. Consequently,
$R_{\sigma LT}(x,Q^{2})=\frac{2Q^{2}}{{\bf q}^{2}}\frac{{\cal R}_{L}({\bf q},\nu)}{{\cal R}_{T}({\bf q},\nu)}.$ (23)
The square of the electric and magnetic form factors for elastic scattering and nuclear excitations are obtained by the integration of the measured response functions over $\nu$ for each nuclear state. It should be noted that when form factors for nuclear excitations are extracted from electron scattering data an additional factor of $Z^{2}$ has traditionally been included in the definition of $\sigma_{M}$ (where $Z$ is the atomic number of the nucleus). However, to keep consistency between the treatment of QE scattering and nuclear excitations we chose not to include the $Z^{2}$ factor in the definition of $\sigma_{M}$ in this analysis.
Figure 6: Top: ${\rm{}^{12}C}$ photo-absorption cross section as a function of photon energy $\nu$ from 0 to 0.6 GeV. Bottom: The response function ${\cal R}_{T}({Q^{2}=0},\nu)$ extracted from photo-absorption cross sections using eq. 24.
II.3.1 Extraction of response functions from photo-absorption cross sections
At $Q^{2}=0$ the virtual photo-absorption cross section is equal to the photo-absorbtion cross section for real photons. Therefore
${\cal R}_{T}(Q^{2}=0,\nu)$ $=$ $\frac{\nu}{2\pi^{2}\alpha}\sigma_{\gamma}(\nu),$ (24)
where $\sigma_{\gamma}(\nu)$ is the photo-absorption cross section for real photons of energy $\nu$.
As real photons can only have transverse polarization, there are no longitudinal photons at $Q^{2}=0$. Therefore, ${\cal R}_{L}(Q^{2}=0,\nu)$ cannot be derived from photo-absorption cross sections and can only be extracted from the universal fit to the electron scattering data,
${\cal R}_{L}(Q^{2}=0,\nu)$ $=$ $\frac{\nu^{3}}{4\pi^{2}\alpha}\lim_{Q^{2}\to 0}\frac{\sigma_{L}(Q^{2})}{Q^{2}}.$ (25)
However, the value of ${\cal R}_{L}(Q^{2}=0,\nu)$ in this region is not relevant since the longitudinal photo-absorption cross section $\sigma_{L}(Q^{2}=0)$ is zero and gives no contribution to the electron scattering cross section.
The measured photo-absorption cross section Bezić et al. (1969); Ahrens et al. (1975); Carrasco and Oset (1992); Bianchi et al. (1996); Muccifora et al. (1999) on 12C is shown in the top panel of Fig 6, while ${\cal R}_{T}(Q^{2}=0,\nu)$ extracted via Eq 24 is shown in the bottom panel. In electron scattering $Q^{2}$=${\bf q}^{2}+\nu^{2}$ must be $\geq 0$. Therefore, for a fixed $\bf q$ the maximum $\nu$ ($\nu^{max}$) occurs when $\nu=\bf q$.
${\cal R}_{T}({\bf q,}\nu=\bf q)$ $=$ $\frac{\bf q}{2\pi^{2}\alpha}\sigma_{\gamma}(\bf q).$ (26)
In photo-absorption the cross section is dominated by the Giant Dipole Resonance, quasi-deuterons Tavares and Terranova (1992) and pion production processes (the cross section for the quasielastic process is zero).
II.4 Experimental extraction of response functions
We use a form of the Rosenbluth method introduced by Jourdan Jourdan (1996, 1995) to separate ${\cal R}_{L}({Q^{2}},\nu)$ and ${\cal R}_{T}({Q^{2}},\nu)$ as follows:
$\Sigma({Q^{2}},\nu)$ $=$ $H\frac{d\sigma}{d\nu d\Omega}$
$=$ $\epsilon{\cal R}_{L}({Q^{2}},\nu)+\frac{1}{2}\left(\frac{{\bf q}}{Q}\right)^{2}{\cal R}_{T}({Q^{2}},\nu),$
with
$H$ $=$ $\frac{1}{\sigma_{M}}{\epsilon}\left(\frac{{\bf q}}{Q}\right)^{4}$ (29)
$=$ $\frac{{\bf q}^{4}}{4\alpha^{2}E^{{}^{\prime}2}}\frac{1}{\cos^{2}(\theta/2)+2({{\bf q}}/{Q})^{2}\sin^{2}(\theta/2).}$
The quantity $\Sigma({Q^{2}},\nu)$ is plotted as a function of the virtual photon polarization $\epsilon$, which varies from 0 to 1 as the scattering angle $\theta$ ranges from 180 to 0 degrees. We use Eq. (29) for $H$ because it is valid for all scattering angles, including $180^{\circ}$. Here, ${\cal R}_{L}({Q^{2}},\nu)$ is the slope, and $\frac{1}{2}\frac{{\bf q}^{2}}{Q^{2}}{\cal R}_{T}({Q^{2}},\nu)$ is the intercept of the linear fit.
However, when Coulomb corrections are included, the above expressions are modified and all the parameters are replaced with “effective parameters” as described below.
II.5 Electron scattering at 180
Including Coulomb corrections the response function ${\cal R}_{T}({Q_{eff}^{2}},\nu)$ can be extracted directly from the electron scattering cross section at 180 using the following expression:
${\cal R}_{T}({Q_{eff}^{2}},\nu)$ $=$ $\left(\frac{E_{0}}{E_{0,eff}}\right)^{2}\frac{{Q_{eff}}^{4}}{4\alpha^{2}E_{eff}^{{}^{\prime}2}}\frac{d\sigma}{d\nu d\Omega},$ (30)
where the $eff$ subscript denotes Coulomb corrected quantities described below.
II.6 ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region
As mentioned earlier, in the nuclear excitation region $4016<E_{x}<$ 40 MeV we use the precise (±$\pm 3\%$) Yamaguchi measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$ (where available). For $E_{x}<16$ MeV we extract ${\cal R}_{L}$ and ${\cal R}_{T}$ from fits Bodek and Christy (2023) to the nuclear excitation form factors, and performing a Gaussian smearing with the width determined by the experimental resolution. At $Q^{2}=0.01$ GeV2 we extract ${\cal R}_{T}$ from electron scattering cross sections at $E_{0}=55$ and 70 MeV and $\theta=180^{\circ}$ published in Goldemberg:64 Goldemberg and Barber (1964) and electron scattering cross sections at $E_{0}=65$ MeV and $\theta$ =180 published in Deforest:65 Deforest et al. (1965). At $Q^{2}=0$ we extract ${\cal R}_{T}$ from measured photo-absorption cross sections.
III Coulomb Corrections
Coulomb corrections to QE and inelastic pion production processes are taken into account using the “Effective Momentum Approximation”(EMA) Aste et al. (2005); Gueye et al. (1999). The approximation is a simple energy gain/loss method, using a slightly higher incident and scattered electron energies at the vertex than measured in the lab. The effective incident energy is $E_{eff}=E_{0}+V_{eff}$, and the effective scattered energy is $E^{\prime}_{eff}=E^{\prime}+V_{eff}$.
Assuming a uniform spherical charge distribution in the nucleus (of radius R) the electrostatic potential inside the charged sphere can be defined as:
$V(r)=\frac{3Z\alpha}{2R}-\frac{Z\alpha}{2R}\left(\frac{r}{R}\right)^{2},$ (31)
where $R$ (in units of GeV) is given by
$R=1.1A^{1/3}+0.86A^{-1/3},$ (32)
and the average potential energy is
$V_{eff}=\frac{4}{5}V(r=0)=\frac{6Z\alpha}{5R},$ (33)
where $Z$ and $A$ are the atomic and mass numbers, respectively.
Figure 7: Comparison to previous ${\cal R}_{L}$ ${\cal R}_{T}$ extractions: (a) Comparisons between our extraction of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ for ${\rm{}^{12}C}$ in units of GeV-1 (red diamonds) to previous extractions by Barreau:82 Barreau et al. (1983a), Jourdan:96 Jourdan (1996, 1995), and Buki:21 Buki and Timchenko (2021) for q=0.3, 0.38 and 0.57 GeV. (b) Comparison of our extraction at $Q^{2}$=0.093 GeV2 to previous extraction by Baran:1988 Baran et al. (1988) at $Q^{2}$=0.1 GeV2 (light green circles) in the region of the $\Delta$(1238) resonance.
However, in this analysis we use a better determination of the effective potential extracted from a comparison of positron and electron QE scattering cross sections on carbon Gueye et al. (1999). The experimentally determined $V_{eff}$ for carbon is $V_{eff}$= 3.1 $\pm$ 0.25 MeV.
Including Coulomb corrections we define:
$E_{0,eff}$ $=$ $E_{0}+V_{eff}$ (34)
$E^{\prime}_{eff}$ $=$ $E^{\prime}+V_{eff}$
$\nu_{eff}=$ $\nu$
$Q^{2}_{eff}$ $=$ $4(E_{0}+V_{eff})(E^{\prime}+V_{eff})\sin^{2}(\theta/2)$
${\bf q_{eff}}^{2}$ $=$ $Q^{2}_{eff}+\nu^{2}$
$W^{2}_{eff}$ $=$ $M^{2}+2M\nu-Q^{2}_{eff}$
The response functions are calculated with $Q^{2}=Q^{2}_{eff}$ and $E^{\prime}=E^{\prime}_{eff}$. In addition, there is a focusing factor $F_{foc}^{2}=\big{[}\frac{E_{0}+V_{eff}}{E_{0}}\big{]}^{2}$ which modifies the Mott cross section. The modified Mott cross section is
$\sigma_{M-eff}$ $=$ $F_{foc}^{2}\frac{\alpha^{2}\cos^{2}(\theta/2)}{[2E_{eff}\sin^{2}(\theta/2)]^{2}}$
Therefore
$\epsilon^{CC}$ $=$ $\left[1+2\left(1+\frac{\nu^{2}}{Q^{2}_{eff}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
$\Sigma^{\prime}({Q^{2}_{eff}},\nu)=H^{CC}\frac{d\sigma}{d\nu d\Omega}$
$=$ $\epsilon^{CC}{\cal R}_{L}({Q_{eff}^{2}},\nu)+\frac{1}{2}\left(\frac{{\bf q_{eff}}}{Q_{eff}}\right)^{2}{\cal R}_{T}({Q_{eff}^{2}},\nu)$
$H^{CC}$ $=$ $\left(\frac{E_{0}}{E_{0}+V_{eff}}\right])^{2}\times$
$\frac{{\bf q_{eff}}^{4}}{4\alpha^{2}E_{eff}^{{}^{\prime}2}}\frac{1}{\cos^{2}(\theta/2)+2({{\bf q_{eff}}}/{Q_{eff}})^{2}\sin^{2}(\theta/2)}$
IV Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$
From the Christy-Bodek overall fit we extract the relative normalization ($N_{i}$) of the various data sets as given in Table 2. For low excitation energy ($E_{x}<$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of excitation energy as described below. Similarly for ($E_{x}>$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of the square of the final state invariant mass $W^{2}$.
We use the fit to apply $Q^{2}$ bin centering corrections ($C_{i}$) to all extracted value off $\Sigma_{i}^{\prime}$ that are within the $Q_{eff}^{2}$ range for each $Q_{eff}^{2}$ and of $W_{eff}^{2}$ bin such that they reflect the values at $Q_{enter}^{2}$, $W_{center}^{2}$, and $\epsilon_{center}$ for each cross section measurement,
$C_{i}$ $=$ $\frac{\epsilon{\cal R}^{fit}_{L}({Q_{center}^{2})}+\frac{1}{2}(\frac{{\bf q_{center}}}{Q_{center}})^{2}{\cal R}^{fit}_{T}({Q_{center}^{2}})}{\epsilon{\cal R}^{fit}_{L}({Q_{eff}^{2})}+\frac{1}{2}(\frac{{\bf q_{eff}}}{Q_{eff}})^{2}{\cal R}^{fit}_{T}({Q_{eff}^{2}})}.$
After the application of the bin centering corrections to all extracted values $\Sigma^{\prime}_{i}$ in each $Q_{eff}^{2}$ bin we can assume that extracted values of $\Sigma_{i}$ are at $Q^{2}_{center}$. When we apply these corrections, we keep the final state invariant mass $W^{2}$ fixed for $\nu>$ 50 MeV, and only correct for the change in $Q_{eff}^{2}$. For $\nu<$ 50 MeV we use the excitation energy $E_{x}$ instead of $W^{2}$.
For $E_{x}$ bins at fixed values of $Q^{2}$ the value of $\nu$ in the bin center is extracted from value of $E_{x}$ in the bin center using the following expression.
$\nu$ $=$ $E_{x}+\frac{Q^{2}}{2M_{A}},$ (36)
where $M_{A}$ is the mass of the nuclear target (for carbon $M_{A}$=11.178 GeV). For $W^{2}$ bins at fixed values of $Q^{2}$, the value of $\nu$ at the bin center is extracted from value of $W^{2}$ in the bin center using the following expression:
$W^{2}$ $=$ $M^{2}+2M\nu-Q^{2},$ (37)
For $E_{x}$ bins at fixed values of $\bf q$ the value of $\nu$ at the bin center is extracted from value of $E_{x}$ in the bin center by solving the following quadratic equation:
$\nu$ $=$ $E_{x}+({\bf q}^{2}+\nu^{2})/(2M_{A})$ (38)
For $W^{2}$ bins at fixed values of $\bf q$ the value of $\nu$ in the bin center is extracted from value of $W^{2}$ in the bin center by solving the following quadratic equation:
$W^{2}$ $=$ $M^{2}+2M\nu-({\bf q}^{2}+\nu^{2}).$ (39)
IV.1 Comparison to Previous Measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
A comparison between our extraction of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ (red diamonds) to previous extractions are shown in Fig. 7 for the three ${\bf q}$ values used in Jourdan:96 Jourdan (1996, 1995). In the Barreau:82 Barreau et al. (1983a) analysis the response functions were extracted for only three $\bf q$ values (0.30, 0.40 and 0.55 GeV) using only their own (Saclay) cross section measurements. A later extraction by Jourdan:96 was also performed for three slightly different values of momentum transfer (${\bf q}=0.30$, 0.38 and 0.57 GeV). In the Jourdan analysis the same Barreau:82 cross sections used but with additional cross section data from SLAC.
Our extraction is more extensive because our data sample is much larger than what was available in 1996 (as summarized in Table 2). In Figure 7(A) we correct the Barreau:82 extractions for the small difference between their $\bf q$ values of 0.40 and 0.55 GeV and the $\bf q$ values of 0.38 and 0.57 GeV, respectively, of Jourdan. In general, our ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are in better agreement with the Jourdan analysis especially at $\bf q$ of 0.57 GeV. In all the plots the universal fit for the total (from all processes) ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ is the solid black line, and the QE component (including Transverse Enhancement) of the fit is the dashed line. A comparison to a recent extraction of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ at ${\bf q}=0.30$ GeV published in Buki:21 Buki and Timchenko (2021) (blue circles) using cross sections measured at Kharkov is also shown. Note that the more recent Kharkov analysis has larger errors and was performed only for ${\bf q}=0.3$ GeV.
There is good agreement between our extraction at $Q^{2}$=0.093 GeV2222 to the extraction by Baran:1988 Baran et al. (1988) at $Q^{2}$=0.1 GeV2 (light green circles) in the region of the $\Delta$(1238) resonance as seen in Figure 7(B).
V Comparison to theoretical predictions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
The extracted ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ (in units of GeV-1) are shown as the red diamonds in Fig. 8-10 for 12 values of $\bf q$. Also shown (where available) are the Yamaguchi:1971 Yamaguchi et al. (1971) measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ in the nuclear excitation region, and ${\cal R}_{T}({\bf q}=0.01)$ GeV extracted from cross sections at 180 (Goldemberg:64 Goldemberg and Barber (1964) and Deforest:65 Deforest et al. (1965)). The values of ${\cal R}_{T}(\bf q,\nu)$ at $\nu$ = $\bf q$ ($Q^{2}$ = 0) are extracted from photo-absorption measurements.
${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ extracted from the Christy-Bodek universal fit (which includes nuclear excitations) are shown as solid black lines, and the QE contributions to the fit (including TE) are the dashed black lines.
Figure 8: Comparison to ED-RMF: Our extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ for ${\rm{}^{12}C}$ for $\bf q$ values of 0.10, 0.148, 0.167, 0.205, and 0.240 GeV versus $\nu$. In the nuclear excitation region we also show measurements from Yamaguchi:1971 Yamaguchi et al. (1971), and ${\cal R}_{T}({\bf q}=0.01)$ GeV extracted from cross sections at 180 (Goldemberg:64 Goldemberg and Barber (1964) and Deforest:65 Deforest et al. (1965)). The values of ${\cal R}_{T}({\bf q}=\nu)$ are extracted from photo-absorption data. In all the plots the universal fit for the total (from all processes) ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ is the solid black line and the QE component (including Transverse Enhancement) of the universal fit is the dotted line. The blue lines are the predictions of the ED-RMF Franco-Munoz et al. (2022, 2023) calculations (which include nuclear excitations).
Figure 9: Comparison to ED-RMF, STA-QMC and GFMC: Same as Fig. 8 for ${\bf q}$ values of 0.30, 0.38, 0.475, 0.57, and 0.649 GeV versus $\nu$. the blue lines are the predictions of ED-RMF, the light green lines are the predictions of STA-QMC and the pink lines are the prediction of GFMC.
Figure 10: Comparison to ED-RMF and STA-QMC: Same as Fig. 8 for $\bf q$ values of 0.756 and 0.991 GeV versus $\nu$. The blue lines are the predictions of ED-RMF, and the light green lines are the predictions of STA-QMC (for $\bf q$=0.756 GeV only). Although STA-QMC calculations have been provided for $\bf q$=0.76 GeV it is expected that that they will be less accurate in this region where relativistic effects, currently not fully included in STA-QMC, become increasingly more relevant.
The dot-dashed vertical orange line is $W=0.93$ GeV, the dot-dashed vertical light green line is $W=1.07$ GeV, the dot-dashed vertical dark blue line is $W=1.23$ GeV and the dot-dashed vertical brown line is where $\nu={\bf q}$ ($Q^{2}=0$ photo-absorption).
It is noteworthy to mention that:
  1. 1.
    The $\Delta$(1232) nucleon resonance is not visible in ${\cal R}_{L}(\bf q,\nu)$ plots and is only seen in ${\cal R}_{T}(\bf q,\nu)$   because the excitation of the $\Delta$ is primarily transverse.
  2. 2.
    At low $\nu$ and low $\bf q$ (or $Q^{2}$) the contributions of nuclear excitation cross sections are significant.
  3. 3.
    The QE peak is not visible at high values of $\bf q$ (or $Q^{2}$) and the cross section is dominated by Fermi motion smeared inelastic pion production processes.
  4. 4.
    At low $\bf q$ and $Q^{2}=0$ ($\nu=\bf q$) the quasielastic cross section is zero and the cross section is dominated by the Giant Dipole Resonance, quasi-deuterons Tavares and Terranova (1992) and pion production processes. The predictions for ${\cal R}_{T}(\bf q,\nu)$ $Q^{2}=0$ are zero for models which do not include these processes.
In the subsections that follow we compare our measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ to the following theoretical predictions:
  1. 1.
    “Green’s Function Monte Carlo” (GFMC)Lovato et al. (2016) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states. GFMC predictions are only available for $0.3\leq{\bf q}\leq 0.57$ GeV.
  2. 2.
    “Energy Dependent-Relativistic Mean Field” (ED-RMF)Franco-Munoz et al. (2022, 2023) which models the contribution of 1-body and two body-currents to the response functions for the single nucleon (1p1h) final state only, but includes nuclear excitations which decay to a single nucleon final state (for both electron and neutrino processes).
  3. 3.
    An improved superscaling model (SuSAv2)Gonzaléz-Jiménez et al. (2014); Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) which uses relativistic mean field to model response functions for the single nucleon (1p1h) final state, and a separate Relativistic Fermi Gas (RFG) modelRuiz Simo et al. (2017) for two nucleon (2p2h) final states (for both electron and neutrino processes).
  4. 4.
    “Short Time Approximation Quantum Monte Carlo” (STA-QMC)Andreoli et al. (2024) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states (for both electron and neutrino scattering processes). Currently, the STA-QMC predictions are only valid for $0.3\leq{\bf q}\leq 0.76$ GeV.
  5. 5.
    “Correlated Fermi Gas” (CFG)Bhattacharya et al. (2024) which models response functions for single nucleon (1p1h) and two nucleon (2p2h) final states.
  6. 6.
    The electron scattering mode of NuWroNuW (2024) which for the electron scattering mode models only the single nucleon final state. Two nucleon final states are included in the neutrino scattering mode.
  7. 7.
    achillesIsaacson et al. (2023) A CHIcago Land Lepton Event Simulator which models the contribution of 1-body and two body-currents to the response functions for the single nucleon final state only (for both electron and neutrino scattering processes). Currently, the achilles  predictions are only valid for $\bf q>$0.5 GeV.
Among these theoretical predictions for $\rm{}^{12}C$ NuWro  is currently used in neutrino experiments, and ED-RMF has recently been included in an update of the NEUT neutrino MC generator. Therefore, in addition to comparing these theoretical predictions to the data, we also include a more extended description of these two theoretical approaches.
The improved SuSAv2 predictions for the 1p1h and 2p2h channels has been implemented in genie  for electrons and neutrinos Dolan et al. (2020); Papadopoulou et al. (2021), and implementation in neut is underway. In addition, the SuSAv2-inelastic model is currently being implemented in genie for electrons and neutrinos. The SuSaA2 2p2h channel has been implemented (only for neutrinos) in NuWro,  andSTA-QMC been implementedBarrow et al. (2021) in genie  but only for $\rm He^{4}$ (though an hadron tensor interfaceGardiner (2019)), and the achilles generator is being developed for use in future neutrino experiments.
V.1 Comparison to the predictions of GFMC, STA-QMC, and ED-RMF
The extracted ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are compared to predictions of GFMC in Fig. 9, STA-QMC in Fig. 9-10) and ED-RMF in Fig. 8-10). The three theoretical calculations include contributions from both 1-body and 2-body (1b+2b) currents. In these three calculations there is enhancement of the calculated transverse response function ${\cal R}_{T}(\bf q,\nu)$ The transverse enhancement (TE) primarily originates from the interference Simons et al. (2022) between 1b and 2b currents which result in 1p1h. (1 nucleon 1 hole) final states. The contribution of 2b currents to ${\cal R}_{T}(\bf q,\nu)$ which results in 2p2h (2 nucleons 2 holes) final states are not included ED-RMF but are included in GFMC and STA-QMC. Note that final states with pions are not included in any of the three calculations (pion production is only included in the improved version of SuSAv20).
The GFMC predictions are shown as the pink lines in Fig. 9. These calculations are rather complex and require a large number of CPU hours, especially at low $\bf q$. Consequently the response functions below $\bf q$=0.3 GeV and above $\bf q$=0.57 are not available for GFMC.
The predictions for $\rm{}^{12}C$ (shown as the light green lines in Fig. 9-10) are provided for $0.3\leq{\bf q}\leq 0.76$ GeV. By its nature STA-QMC only works for $\bf q>$ 0.3 GeV. The STA-QMC predictions for for $0.3\leq{\bf q}\leq 0.6$ GeV are in agreement with our measurements. Presently, STA-QMC does not fully account for relativistic effects at the vertex. Therefore, although STA-QMC calculations have been provided for $\bf q$=0.76 GeV (shown in Fig. 10), it is expected that that they will be less accurate in the high $\bf q$ region where relativistic effects become increasingly more relevant. Although the STA-QMC calculations are based on the same many-nucleon Hamiltonian and electromagnetic currents as GFMC, the STA-QMC calculations are in better agreement with our measurements than GFMC.
The ED-RMF theoretical predictions are shown as the blue lines. These predictions (which are available at all values of $\bf q$) include contributions to 1p1h final states from QE scattering as well as from nuclear excitations. At higher values of $\bf q$ and higher values of $\nu$ the ED-RMF predictions for ${\cal R}_{T}(\bf q,\nu)$ are below the data. This is expected because the ED-RMF theoretical calculations for ${\cal R}_{T}(\bf q,\nu)$ do not include 2-body current processes with 2p2h final states or Fermi smeared pion production processes. The STA-QMC calculation includes both single and two nucleon final states but is only valid over a more restricted kinematic range and does not account for nuclear excitations. The calculations of both ED-RMF and STA-QMC are directly applicable to the same kinematic regions for neutrino scattering.
V.1.1 The ED-RMF theoretical formalism
Because the ED-RMF approach has recently been implemented in the NEUT neutrino generator, we include additional details on the ED-RMF approach. In the ED-RMF approach the hadronic current operator includes one- and two-body current contributions (see details in ref. Franco-Munoz et al. (2022)). In this 1p1h calculation, only the contributions from transitions where one nucleon below the Fermi level (hole) is promoted to a continuum state above the Fermi level (particle) by the current operator are included. Matrix elements of this current operator in between hole-particle combinations are computed in an unfactorized fashion to compute the cross-section. The momentum distribution of the bound nucleons in the initial state are obtained by solving the Diract equation with the Relativistic Mean Field as in ref. Sharma et al. (1993). The occupations and energies are chosen to be consistent with a representation of a semi-phenomenological spectral function similar to the Rome spectral function for ${}^{12}C$. A continuous missing-energy profile and occupations is employed as in ref. Franco-Patino et al. (2022), i.e. 3.3 nucleons for the 1p3/2 and 1.8 nucleons for the 1s1/2 shell for both protons and neutrons. The remaining 0.8 protons and neutrons, are ascribed to the non-discrete content of the spectral function, represented by a high missing-energy and momentum contribution.
For the final state, the momentum distributions for the particle states in the continuum are computed by solving the Dirac equation with the energy-dependent relativistic mean-field (ED-RMF) potential, which has no imaginary part so no flux is lost. The ED-RMF is the same potential used in the bound state but multiplied by a function that weakens it for increasing energy (for see details ref. González-Jiménez et al. (2019); Nikolakopoulos et al. (2019)). At low energies of the final nucleon (up to about 20 MeV in the continuum) the ED-RMF potential is the same potential employed to compute the momentum distributions for the initial bound nucleons an at larger energies it approaches the behavior of a phenomenological optical potentials. This ensures orthogonality between the initial and final states at these low energies and thus the reliability of the cross-section computed in this region. In this way, the ED-RMF calculation also preserves the contribution (and resonant character) from single-channel resonances corresponding to excitations appearing slightly above the single-nucleon emission threshold of the 12C final system, which subsequently decay by single nucleon emission.
V.1.2 Comparisons to ED-RMF in the nuclear excitation region
Figure 11: Nuclear excitation Region for fixed q bins vs $E_{x}$: Comparison of measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ in the nuclear excitation region to the predictions of ED-RMF ( blue lines), GFMC ( pink lines) and STA-QMC ( light green lines). The GFMC and STA-QMC predictions are only available for q values of 0.3, 0.38 and 0.57 GeV.
As seen in Fig. Fig. 8 - Fig. 9 the contribution of nuclear excitation to the response functions is significant for $\bf q$ values less than 0.3 GeV. In this region, for excitation energies above 20 MeV the ED-RMF predictions are in reasonable agreement (on average) with the measurements. Fig. 11 is an expanded view of the nuclear excitation region, where the excitations of the nuclear targets can be seen above the QE contribution (the dashed black line). Here ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are shown versus excitation energy $E_{x}$. The ED-RMF calculations require one nucleon in the continuum in the final state, thus they show strength only above the one-nucleon separation energy (16 MeV for a proton and 18.7 MeV for a neutron in $\rm{}^{12}C$).
Above the one-nucleon separation threshold and small energy transfers, where the nucleon in the continuum has less than 10 MeV, typically, the ED-RMF calculation exhibits single-channel resonances of the RMF potential employed to compute the continuum wave functions for the nucleon in the final state. At these low energies of the nucleon, one expect these resonances of the potential to correspond to actual resonances of the target nucleus, with the correct cross sections, if the potential is realistic enough, in a phenomenological sense. The ED-RMF calculations are performed with some moderate binning in energy, which contributes to the averaged smoothed look of the ED-RMF predictions. As seen in Fig. 11 for $E_{x}$ above the nucleon separation energy the ED-RMF predictions in the nuclear excitation region for $\bf q$ values less than 0.3 GeV are reasonable (on average). This indicates that the ED-RMF potential used is a reasonable approximation of the actual nuclear potential. Since the same potential is used in the ED-RMF calculations of the response functions in neutrino scattering, the resulting response functions in the nuclear excitation region for neutrinos would probably also be a reasonable approximation to the actual nuclear excitations in neutrino scattering. None of the other models used in neutrino MC generators to date include the contribution of nuclear excitations.
As mentioned earlier, the ED-RMF response functions incorporate the average contribution from nuclear excitations which decay via single nucleon emission. Decay modes which decay to other channels would not, however, contribute to ED-RMF responses. For example, decays with $\alpha$ particles, $\gamma$’s or deuterons in the final state are not included. The ED-RMF predictions in the nuclear excitation region above 20 MeV are expected to be reasonable because the predominant decay modes of the nuclear excitations are to single nucleon final states.
For example, the relative branching ratios for the decays of nuclear excitations between 12 and 21 MeV to protons and $\alpha$ particles have been measuredNeuschaefer and Tabor (1985). The decays of the 12.71 and 16.11 MeV excitations (which are below the proton and neutron separations energy) are primarily via the the emission of $\alpha$ particles, while for the 20.62 MeV excitation the branching ratio to decay via proton emission is a factor of 2 higher than the decay via $\alpha$ emission. Note that Christy-Bodek fit to inclusive electron scattering data describes all nuclear excitations irrespective of their decay modes.
Figure 12: Comparison to CFG, and NuWro for fixed q bins: Our extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ for $\bf q$ values of 0.10, 0.148, 0.167, 0.205, and 0.240 GeV versus $\nu$. In the nuclear excitation region we also show measurements from Yamaguchi:1971 Yamaguchi et al. (1971), and ${\cal R}_{T}({\bf q}=0.01)$ GeV extracted from cross sections at 180 (Goldemberg:64 Goldemberg and Barber (1964) and Deforest:65 Deforest et al. (1965)). The values of ${\cal R}_{T}({\bf q}=\nu)$ are extracted from photo-absorption data. The universal fit for the total (from all processes) ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ is the solid black line and the QE component (including Transverse Enhancement) of the universal fit is the dotted line. The solid brown line is the prediction of the “Correlated Fermi Gas” (CFG) Bhattacharya et al. (2024) calculation.The solid pink line is the prediction of the NuWro Monte Carlo Generator using a spectral function formalism (labeled NuWro-SF) , and the solid green line is the predictions of NuWro-SF and accounting for the effect of final state interaction (labeled NuWro-SF-FSI). (Note that ${\cal R}_{T}(\bf q,\nu)$ in NuWro is underestimated because the electron scattering mode does not include an MEC model).
Figure 13: Comparison to CFG, NuWro and achilles  for fixed q bins: Same as Fig. 12 for ${\bf q}$ values of 0.30, 0.38, 0.475, 0.57, and 0.649 GeV versus $\nu$. The the solid blue squares are the predictions of achilles  (for $\bf q$ = 0.57 and 0.659 GeV). (Note that ${\cal R}_{T}(\bf q,\nu)$ in NuWro is underestimated because the electron scattering mode does not include an MEC model).
Figure 14: Comparison to CFG, NuWro and ACHILLESS for fixed q bins: Same as Fig. 12 for $\bf q$ values of 0.756, 0.991, 1.619 1.92, and 2.213 GeV versus $\nu$. The the solid blue squares are the predictions of achilles  (for $\bf q$ = 0.756, 0.991 and 1.619 GeV).(Note that ${\cal R}_{T}(\bf q,\nu)$ in NuWro is underestimated because the electron scattering mode does not include an MEC model).
Figure 15: Comparison to NuWro for fixed q bins: Same as Fig. 12 for $\bf q$ values of $\bf q$ values of 2.5, 2.763, and 3.5 GeV versus $\nu$.
V.2 Comparison to NuWro, CFG and ACHILLES
In this section we compare the extracted ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$  for 18 fixed values of $\bf q$ (shown in Fig.12 - 15) to the predictions of the NuWro-SF (spectral function) calculation and the predictions of NuWro-SF-FSI (spectral function calculation with Final State Interaction). In addition, we compare the data to the predictions of the Correlated Fermi Gas (CFG), and to the predictions of achilles (which are only valid for $\bf q>$0.5 GeV). Currently, the NuWro generator is being used in neutrino experiments, and achilles is being developed for use in neutrino experiments. Therefore, these two calculations also described in more detail below.
V.2.1 The NuWro Monte Carlo Generator
The NuWro Monte Carlo generator NuW (2024) is designed to simulate neutrino-nucleus interactions in the few-GeV energy region, relevant for accelerator-based experiments. It has been primarily developed by the neutrino-theory group at the University of Wrocław since 2004 Sobczyk et al. (2005); Juszczak et al. (2005).
To model lepton interactions with atomic nuclei, NuWro assumes the validity of the impulse approximation, in which the process of scattering is assumed to involve predominantly a single nucleon, with the remaining nucleons acting as a spectator system.
Simulating neutrino interactions with nucleons, NuWro accounts for charged-current quasielastic (QE) scattering Llewellyn Smith (1972), hyperon production Thorpe et al. (2021), single pion production Sobczyk et al. (2005), and deep-inelastic scattering Bodek and Yang (2002). In interactions with atomic nuclei, NuWro also takes into account mechanisms involving two-body currents Bonus et al. (2020) and coherent pion production Berger and Sehgal (2009).
In the electron mode, NuWro currently simulates only QE scattering Banerjee et al. (2024), with several options available to model nuclear effects. The results presented in this article are obtained within the spectral function (SF) approach Benhar et al. (2005) (without and with FSI).
NuWro uses the carbon SF of Ref. Benhar et al. (1994), which consistently combines the shell structure determined in coincidence electron-scattering experiments at Saclay Mougey et al. (1976) with the results of theoretical calculations for infinite nuclear matter at different densities Benhar et al. (1989), by employing the local-density approximation.
Happening inside the nucleus, the interaction between a lepton and a nucleon is affected by the spectator system. The surrounding nucleons interact with the struck nucleon, which modifies its energy spectrum and leads to a more complicated energy conservation in the vertex than for a free nucleon. These effects of final-state interactions (FSI) induce a shift and a broadening of the double differential cross section in a kinematics-dependent manner. In NuWro, they are implemented following Ref. Ankowski et al. (2015). In addition, the effect of Pauli blocking is accounted for by the action of the step function Benhar et al. (2005), with the average Fermi momentum $\overline{p}_{F}=211$ MeV Ankowski et al. (2015).
In the context of the ${}^{12}{\rm C}(e,e^{\prime})$ response functions, it is important to bear in mind that the electron mode of NuWro does not account for mechanisms of interaction other than QE scattering induced by the one-body current. The two-body currents and interference between one-body and 2-body currents are responsible for the enhancement of the transverse response functions. Therefore, in electron scattering mode NuWro is expected to underestimate the transverse response function in the QE region. and the comparisons with NuWro are only valid for ${\cal R}_{L}(\bf q,\nu)$
In addition, in electron scattering mode NuWro does not include pion production processes. Moreover, discrepancies with experimental data are expected to occur for the processes beyond the impulse approximation, such as nuclear excitations and the excitation of the giant dipole resonance, which are not included in the electron scattering or neutrino scattering modes in NuWro.
V.2.2 Comparison to NuWro  predictions
As seen in Fig. 12 and 15 the predictions of NuWro with FSI are in better agreement with the data than the prediction without FSI, especially at low $\bf q$. (The effect of FSI above $\bf q$ of 0.65 GeV is small). Therefore, we focus on comparisons to the predictions of NuWro with FSI (NuWro-SF–FSI, solid solid green curve).
As mentioned above, NuWro in electron scattering mode does not include a model for the enhancement in ${\cal R}_{T}$ from 2-body currents. Therefore, as expected the predictions of NuWro-SF-FSI are lower than the data.
The predictions of NuWro-SF-FSI for ${\cal R}_{L}(\bf q,\nu)$ at low $\bf q$ are higher than the data, since they do not account for the quenching of ${\cal R}_{L}^{QE}(\bf q,\nu)$ at small values of $\bf q$. However, since NuWro-SF-FSI also does not include a model for nuclear excitations, the overestimate of ${\cal R}_{L}(\bf q,\nu)$ partially compensates for the missing nuclear excitations in the longitudinal channel.
The direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek universal fit can be used to extract corrections (e.g. tuning) that will bring the predictions into better agreement with the data as discussed in section 3.
V.3 Comparison to the prediction of the Correlated Fermi Gas (CFG)
The extraction of the Fermi momentum parameters in the Relativistic Fermi Gas (RFG) model from fits to electron scattering cross sections on various nuclei (at $\bf q$=0.45 GeV) were published by Moniz and collaborators in 1971Moniz et al. (1971). In 1981, Bodek and RitchieBodek and Ritchie (1981a, b) added the contribution of two-nucleon correlations to the model. That models was developed for analysis of deep inelastic scattering on nuclear targets at high values of $Q^{2}$.
The recent Correlated Fermi Gas model Bhattacharya et al. (2024) includes two-nucleon correlations in a more sophisticated way. As seen in Fig.12 - 15 the model breaks down at low momentum transfers and works somewhat better at higher momentum transfers. Here also, the direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek universal fit can be used to extract corrections (e.g. tuning) that will bring the predictions into better agreement with the data as discussed in section 3.
V.4 Comparison to the predictions of achilles
The achilles  predictionsSteinberg (2024) for ${\bf q}$ values of 0.570, 0.649, 0.756, 0.991, and 1.659 GeV are shown as solid blues squares in Fig.13 - 14..
V.4.1 The achilles approach
The achilles  predictions are computed in the impulse approximation using a quantum Monte Carlo (QMC) based spectral function. The contributions from 1-body currents and the interference between 1-body and 2-body currents, both leading to 1-nucleon knockout (1p1h) are included. These are computed using the $\rm{}^{12}C$ QMC spectral function used in Korover et al. (2023); Simons et al. (2022); Lovato et al. (2023a, b) and are corrected for elastic FSI using a folding function. The interaction mechanisms which are included contribute almost all of the strength in the QE peak and lower in $\nu$. Not included are the contribution from two nucleon knockout by 2-body currents or pion production, which contribute at higher energy transfer. Pauli blocking is implemented as a simple fermi gas prescription, where the cross section is put to 0 if the final state nucleon momentum is less than $k_{F}$ ($k_{F}$ = 225 MeV for $\rm{}^{12}C$ is used). The achilles  predictions are only valid for $\bf q>0.5$ GeV.
V.4.2 Comparison to achilles predictions
As seen in Fig.13 and Fig.14 the achilles  predictions for ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are in better agreement with the data than the predictions of NuWro-SF-FSI. For ${\cal R}_{T}(\bf q,\nu)$ this is expected because achilles  also include the contribution of 2-body currents and the interference between 1-body and 2-body currents leading to 1-nucleon final states. The predictions for ${\cal R}_{T}(\bf q,\nu)$ are lower than the QE+TE component of Christy-Bodek fit because the transverse enhancement resulting in 2-nucleon final states (2p2h) are not included in the model.
Figure 16: Comparison to ED-RMF and NuWro for fixed $Q^{2}$ bins: Same as Fig. 12 for $Q^{2}$ values of 0.01 0.02, 0.026, 0.04 and 0.056 GeV2 versus $\nu$. ( ${\cal R}_{T}(\bf q,\nu)$ in NuWro is underestimated because the electron scattering mode does not include an MEC model).
Figure 17: Comparison to ED-RMF and NuWro for fixed $Q^{2}$ bins: Same as Fig. 12 for $Q^{2}$ values of 0.093 0.12, 0.16, 0.265 and 0.38 GeV2 versus $\nu$. (Note that ${\cal R}_{T}(\bf q,\nu)$ in NuWro is underestimated because the electron scattering mode does not include an MEC model).
Figure 18: Comparison to ED-RMF and NuWro for fixed $Q^{2}$ bins: Same as Fig. 12 for $Q^{2}$ values of 0.50, 0.8, 1.25, 1.75 and 2.25 GeV2 versus $\nu$. (Note that ${\cal R}_{T}(\bf q,\nu)$ in NuWro is underestimated because the electron scattering mode does not include an MEC model).
Figure 19: Comparison to ED-RMF for fixed $Q^{2}$ bins: Same as Fig. 12 for $Q^{2}$ values of 2.75, 3.25, and 3.75 GeV2.
Figure 20: Nuclear excitation Region for fixed $Q^{2}$ bins vs $E_{x}$: Comparison of measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as a function of $E_{x}$ for Q2 values of 0.01, 0.026, 0.056, 0.12 GeV2 to the predictions of ED-RMF.
V.5 Comparisons of ${\cal R}_{L}(\bf q,\nu)$ to the predictions of NuWro-SF-FSI and ED-RMF in 18 fixed $Q^{2}$ bins
In the previous sections we compare the extracted values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ to the predictions of NuWro-SF-FSI and ED-RMF for fixed values of $\bf q$. Comparisons versus $\bf q$ are typically used to validate nuclear theory predictions.
For neutrino scattering the validation of MC generators at fixed values of $Q^{2}$ is more relevant. Therefore, in this section we present the comparison of the extracted values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ to the predictions of both NuWro-SF-FSI and ED-RMF 18 fixed values of $Q^{2}$ as shown in Fig.16 - 19.
As expected, the NuWro predictions for ${\cal R}_{T}(\bf q,\nu)$ are underestimated because the electron mode of NuWro does not include the contribution of Meson Exchange Currents (which are included in the neutrino version of NuWro). Therefore, for NuWro the relevant validations are only for ${\cal R}_{L}(\bf q,\nu)$
At low $Q^{2}$ we find that the NuWro-SF-FSI predictions are in better agreement with the data than the predictions of NuWro-SF. At high $Q^{2}$ we find that there is little difference between the predictions of NuWro-SF-FSI and NuWro-SF. Therefore, we only show the predictions of NuWro-SF-FSI at low $Q^{2}$ and NuWro-SF at high $Q^{2}$.
Fig. 20 is an expanded view of the nuclear excitation region, where the excitations of the nuclear targets can be seen above the QE contribution (the dashed black line). Here the response functions are shown versus excitation energy $E_{x}$. This figure is the same as Fig. 11 except that it is for bins in fixed $Q^{2}$ instead of bins in fixed q.
As seen in Fig.16 - 20 the ED-RMF predictions for ${\cal R}_{L}(\bf q,\nu)$ are in much better agreement with experiment than the NuWro-SF-FSI predictions especially at low $Q^{2}$ where the contribution from nuclear excitations is significant.
VI Comparison to the improved superscaling model SuSAv2
The improved superscaling model (SuSAv2) is described in detail in references Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023). Among the models that we have investigated, the recently improved SuSAv2 is the only model that includes QE, 2p2h-MEC, and pion production processes including resonance production, Shallow Inelastic Scattering (SIS) and Deep Inelastic (DIS) processes, (but does not include nuclear excitations). In this publication we focus on the comparison of the SuSAv2 response functions for the QE-1p1h and 2p2h-MEC processes. Investigation of pion production processes will be reported in a future publication.
The SuSAV2 predictions were compared to few electron scattering cross section measurement in references Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023). Here we present the first comparison of the SuSAv2 predictions to measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ spanning the complete kinematic range of $Q^{2}$ and $\nu$ of interest to neutrino experiments. These comparisons are shown for the 18 values of fixed $\bf q$ (Figures 21-24) and the 18 values of fixed $Q^{2}$ (Figures 25-28) in Appendix A.
In all the plots the universal fit for the total (from all processes) ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ is the solid black line and the QE component (including Transverse Enhancement) of the universal fit is the dotted line. The predictions of SuSAv2 are the dashed pink lines for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of ED-RMF Franco-Munoz et al. (2022, 2023) for QE-1p1h (including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC-2p2h model to account for the 2p2h contribution to the response functions (solid green line).
VI.1 SuSAV2 QE(1p1h)
In the original SuSA model, which is based on a fit to the electron longitudinal response functions, the longitudinal and transverse scaling functions are equal, thus resulting in an underestimation in the transverse channel. In the improved SuSAv2 the Relativistic Mean Field (RMF) approach is used to obtain the theoretical longitudinal and transverse scaling functions. The SuSAv2 RMF transverse scaling function is somewhat larger than the RMF longitudinal scaling function, thus resulting in a transverse enhancement of the QE-1p1h cross section. However differences between SuSAv2-QE-1p1h and ED-RMF-QE-1p1h are expected at small q, $\nu$ and $Q^{2}$ where scaling violations present in the RMF theory (and also nuclear excitations) as well as the effect of 2-body currents leading to 1p1h in the ED-RMF approach are not fully incorporated in the SuSAv2 approach.
We find that SuSAV2-QE-1p1h predictions for ${\cal R}_{L}^{QE}(\bf q,\nu)$ (dashed pink lines) are somewhat higher than our measurements (especially at small q), and the predictions for ${\cal R}_{T}^{QE}(\bf q,\nu)$ are somewhat lower than our measurements (especially at small q). Although ED-RMF and SuSAv2 cross sections are similar apart from the very low-energy region, some important differences appear when studying the ${\cal R}_{L}^{QE}(\bf q,\nu)$ and ${\cal R}_{T}^{QE}(\bf q,\nu)$ channels separately. These differences can be mainly ascribed to the lack of 2-body currents leading to 1p1h in the SusAv2 approach. When comparing theory predictions to experimental electron scattering cross sections, such differences may be hard to observe because an overestimation of ${\cal R}_{L}^{QE}(\bf q,\nu)$ can be partially compensated by an underestimation of ${\cal R}_{T}^{QE}(\bf q,\nu)$ unless the comparisons include both cross section data at very small angles and cross section data at very large angles.
A more significant issue is that for low $\bf q$ and low $Q^{2}$ ($\bf q<$ 0.3 GeV and $Q^{2}<$ 0.1 GeV2) the SuSAv2-QE-1p1h predictions for ${\cal R}_{L}^{QE}(\bf q,\nu)$ and ${\cal R}_{T}^{QE}(\bf q,\nu)$ are shifted to low values of $\nu$ which results in unphysical cross sections (at negative excitation energies) at small $\nu=0$.
In summary,, additional corrections and tuning of the SuSAv2-QE-1p1h calculations are needed to achieve better agreement with our measurements.
VI.2 SuSAV2 MEC-2p2h
In the Meson-Exchange-Current 2p2h model (MEC-2p2h) incorporated in SuSAv2 the main contribution in electron scattering is transverse and the contribution in the longitudinal channel is small. This is not the case for neutrinos where the 2p2h longitudinal axial contribution is important. In electron scattering the enhancement in the QE peak region originating from the 1p1h RMF transverse enhancement is in general bigger than the 2p2h contribution. The 2p2h channel becomes more important in the dip region between the QE and $\Delta$(1238) peaks. The sums of the predictions of SuSAv2-QE-1p1h and SuSAv2–MEC-2p2h are shown as the solid pink lines in the plots in Appendix A. In general the inclusion of the 2p2h contribution results in better agreement with our measurements in the dip region. However, at high values of q and $Q^{2}$ (e.g. $\bf q>$1.6 GeV and $Q^{2}>$ 1.25 GeV) the SuSAv2 MEC-2p2h model predictions for the longitudinal response functions have unphysical peaks. This could be because the parameterization of the 2p2h RFG-based nuclear responses from Ruiz-Simo et al.Ruiz Simo et al. (2017) included in this implementation of the SuSAv2-MEC model is valid only for ${\bf q}<2$ GeV and $\nu<2$ GeV (although it can be extended to higher kinematics).
In contrast the SuSAv2-MEC-2p2h implemented in genie  uses interpolation methods and therefore is more accurate than the parameterization formulae used in this implementation. This can introduce important differences at low kinematics (${\bf q}<$ 0.3 GeV) and also at very high kinematics (${\bf q}>$1.5 GeV).
Since the ED-RMF-QE-1p1h predictions are in better agreement with our measurements than the predictions of SuSAv2-QE-1p1h, we also show the sum of the predictions of ED-RMF-QE-1p1h and the predictions of SuSAv2-MEC-2p2h as the solid green lines in the plots in Appendix A. Combining the ED-RMF-QE-1p1h predictions with the SuSAv2-MEC-2p2h predictions provides a better description of the response function for both single nucleon (from QE and nuclear excitations) and two nucleon final states.
Model Currents Final State Validity RL RT high $\bf q$ low $\bf q$
ED-RMF 1b,2b 1p1h, nucl. exci. $all\leavevmode\nobreak\ \bf q$ OK OK 2p2h model 2p2h model
QMC no 2p2h need 2p2h need 2p2h needed needed
model model use SuSAV2 ? use SuSAv2?
SuSAV2 1b 1p1h + $all\leavevmode\nobreak\ \bf q$ low $\nu$ unphysical low $\nu$ unphysical OK need RL quench.
RMF Scaling for 2p2h model low q high low q low need more RT
function 1p1h no nucl exc. need quench
STA-QMC 1b,2b 1p1h, 2p2h $0.650.3<{\bf q}<$ 0.65 OK OK relativistic analytic extrapol.
no nucl exc. only corr. needed needed
NuWro SF-FSI 1b 1p1h all q too high too low needs needs
e-mode spectral no nucl exc. needs needs 2p2h 2p2h
function no 2p2h quenching enhancement model model
achilles 1b,2b 1p1h ${\bf q}>$0.5 OK OK RL RT need other
QMC spectral no 2p2h. only need need need low q models
function no nucl exc. 2p2h model 2p2h model 2p2h model for 1p1h, 2p2h
GFMC 1b,2b 1p1h, 2p2h $0.570.3<{\bf q}<$ 0.57 low $\nu$ unphysical OK CPU intensive CPU intensive
QMC no nucl. exc. only q=0.57 too high not possible not possible
CFG 1b 1p1h,2p2h all q poor poor poor poor
Correlated no nucl. exc. agreement agreement agreement agreement
Fermi Gas
Table 3: A summary of comparisons of $\rm{}^{12}C$ ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ to theoretical predictions (q units are in GeV).
Because our extractions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from all available data electron scattering data span a large range of $\bf q$ and $Q^{2}$, the comparisons can be used extract $\bf q$ dependent corrections to nuclear models and neutrino/electron MC generators to achieve better agreement our measurement s (as has been done in the Christy-Bodek fit). For example, in the QE and nuclear excitation region:
  1. 1.
    The locations of the QE in $\nu$ are sensitive to the removal energy (energy shift) which shifts the QE peak to higher values of $\nu$) and also a $\bf q$ dependent final state optical potential Bodek and Cai (2019) which shifts the QE peak to lower values of $\nu$ at low $\bf q$. Tuning these parameters in SuSAv2 would bring the position of the QE peaks into better agreement with the measurements at low $\bf q$ and $Q^{2}$. In addition, introducing a q-dependence in the RMF-based SuSAv2 scaling functions (which are roughly constant for ${\bf q}>$ 0.3 MeV, but not for smaller values of q) can be used to incorporate into SuSAV2 the the low-energy RMF nuclear effects that break the scaling behavior at low q, which will modify the shape and magnitude of the SuSAv2 QE predictions at that region.
  2. 2.
    The location of the $\Delta$(1232) is also sensitive to removal energy and and final state optical potential. We observe that the shift of the peak of the $\Delta$(1232)  Bodek and Cai (2020) to lower values of $\nu$ is larger than the shift of the QE peak. This is unexpected because the optical potential for the $\Delta$(1232) is about 2/3 of the optical potential for the nucleonEricson and Weise (1988). In SuSAv2 this larger shift is accounted for by the 2p2h-MEC contribution to the cross section in the dip region. In gibuu this shift is accounted for by a nuclear pion production continuumMosel and Gallmeister (2024). In the Christy-Bodek fit this additional shift is parametrized in terms of a larger “effective optical potential” Bodek and Cai (2020) for the $\Delta$(1232) and inelastic pion production processes. Alternatively, the treatment of the real and imaginary parts of the $\Delta$ propagator in 2p2h and resonance channels and $\Delta$ in-medium modifications can also be tuned or modified to achieve better agreement with electron scattering data.
  3. 3.
    The magnitude of ${\cal R}_{L}(\bf q,\nu)$ can tuned be via a $\bf q$ dependent longitudinal quenching factor, and the amplitude of ${\cal R}_{T}(\bf q,\nu)$ can be tuned via an $\bf q$ dependent additive transverse enhancement contribution (as done in the Christy-Bodek fitBodek and Christy (2022)). Including and tuning such parameters in SuSAv2 would bring the magnitude of ${\cal R}_{L}^{QE}(\bf q,\nu)$ and ${\cal R}_{T}^{QE}(\bf q,\nu)$ predictions into better agreement with the measurements.
  4. 4.
    In ED-RMF nuclear excitations (which decay to single nucleons in the final state) occur via the interactions of final state nucleons with the Energy Dependent Mean Field. In the other calculations, a model describing nuclear excitations (with excitation energies above the proton separation energy) can be added.
  5. 5.
    A model describing nuclear excitations (with excitation energies below the proton separation energy) which decay via $\gamma$ ray emission is included in the inclusive Christy-Bodek fit Bodek and Christy (2023). At present none of the theoretical predications include nuclear excitations which decay via $\gamma$ emission.
VII Summary
We have done a global extraction of the ${\rm{}^{12}C}$ longitudinal (${\cal R}_{L}$) and transverse (${\cal R}_{T}$) nuclear electromagnetic response functions from an analysis of all available inclusive electron scattering cross section on ${}^{12}C$. The response functions are extracted for a large range of energy transfer $\nu$, spanning the nuclear excitation, quasielastic, resonance and inelastic continuum and over a large range of $Q^{2}$. We extract ${\cal R}_{L}$ and ${\cal R}_{T}$ as a function of $\nu$ for fixed values of both $Q^{2}$ (0 $\leq Q^{2}\leq 3.5$ GeV2) and $\bf q$ (0.1 $\leq{\bf q}\leq 3.75$ GeV).
Since the extracted response functions cover a large range of $Q^{2}$ and $\nu$, they can be readily used to validate nuclear models as well as validate and tune Monte Carlo generators for electron and neutrino scattering experiments. Alternatively, the Christy-Bodek global fit to all electron scattering data on $\rm{}^{12}C$ can also be used to validate nuclear models and also validate and tune Monte Carlo generators for electron and neutrino scattering experiments over a larger region of $\nu$. In this paper we focus on comparison to models in the nuclear excitation region, and for single nucleon (QE-1p1h) and two nucleon (2p2h) final states.
Because the mass of the muon is much larger than the mass of the electron there are kinematic regions in electron-neutrino charged current scattering which are not accessible in the scattering of muon-neutrinos. Therefore, in neutrino oscillation experiments the modeling of the cross sections at low momentum transfer and low energy transfer is important for accounting for the difference between muon-neutrino and electron-neutrino cross sections. This is the region where the contribution of nuclear excitations is significant. At present, only ED-RMF accounts for nuclear excitations which decay to a single nucleon in the final state (but not excitations which decay via gamma emission).
We have compared measurements in the QE region to the predictions of several theoretical calculations as summarized in Table 3. In particular ED-RMF, STA-QMC and SuSAv2 have the added benefit that the calculations are also directly applicable to the same kinematic regions for neutrino scattering. We find that the SuSAV2-QE-1p1h predictions for ${\cal R}_{L}^{QE}(\bf q,\nu)$ are somewhat higher than our measurements and the predictions for ${\cal R}_{T}^{QE}(\bf q,\nu)$ are a somewhat lower than our measurements. In addition, at small momentum transfers the SuSAV2-QE-1p1h predictions for ${\cal R}_{L}^{QE}(\bf q,\nu)$ and ${\cal R}_{T}^{QE}(\bf q,\nu)$ are shifted to lower values of $\nu$ which result in unphysical cross sections at small $\nu$. These differences can be remedied via the introduction of additional tuning parameters in the theory.
Among all the QE theoretical predictions, ED-RMF provides the best description of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ for QE scattering (for 1p1h final states) over the largest kinematic range in $Q^{2}$ and $\nu$. These predictions (which are available at all values of $\bf q$) include contributions to 1p1h final states from QE scattering as well as from nuclear excitations. At higher values of $\bf q$ and higher values of $\nu$ the ED-RMF predictions for ${\cal R}_{T}(\bf q,\nu)$ are below the data. This is expected because the ED-RMF theoretical calculations for ${\cal R}_{T}(\bf q,\nu)$ do not include processes with two nucleons or pions in the final states. Combining the ED-RMF-QE-1p1h predictions with the SuSAv2-MEC-2p2h predictions provides a better description of both single nucleon (from QE and nuclear excitations) and two nucleon final states. In our next publication (in which the resonance region will be investigated) we plan to include ED-RMF-1-pion production model.
The STA-QMC calculation includes both single and two nucleon final states but at present it is only valid over a more restricted kinematic range and does not account for nuclear excitations. Extension of the theory to higher and lower values of momentum transfer are currently under way.
In future communications we plan to report on the comparisons of response functions (including the nucleon resonance region) for $\rm{}^{12}C$, $\rm{}^{40}Ca$ and $\rm{}^{56}Fe$ to the predictions of ED-RMF, SuSAV2 and the genie MC generator.
VIII Supplemental Materials
The extracted ${\cal R}_{L}$ and ${\cal R}_{T}$ values (shown in the figures) are available as supplemental materials sup . In addition, the values of the Christy-Bodek fit to ${\cal R}_{L}$ and ${\cal R}_{T}$ over a larger range of $\nu$ (and also listing separately the contributions of nuclear excitations, quasielastic scattering, transverse enhancement, and pion production processes) are available as supplemental materials.
VIII.1 Acknowledgments
We thank Noah A. Steinberg for providing the predictions from the achilles  MC generator. We thank Lorenzo Andreoli for providing the predictions from STA-QMC and Saori Pastore for clarification of the predictions of STA-QMC. We thank Sam Carey for providing the predictions from CFG. We thank Tania Franco Munoz for providing the predictions from ED-RMF, and Jose Manuel Udias and Raul Gonzalez-Jimenez for clarification of the predictions of ED-RMF.
Research supported in part by the Office of Science, Office of Nuclear Physics under contract DE-AC05-06OR23177 (Jefferson Lab) and by the U.S. Department of Energy under University of Rochester grant number DE-SC0008475. The work of A. M. Ankowski is partly supported by the Polish National Science Centre under grant UMO-2021/41/B/ST2/02778. The work of Julia Tena Vidal is supported by a European Research Council ERC grant (NeutrinoNuclei, 101078772). G. D. Megias’ work is supported by the Spanish Ministerio de Ciencia, Innovación y Universidades and ERDF (European Regional Development Fund) under contracts PID2020-114687GB-100 and PID2023-146401NB-I00, by the European Union’s Horizon 2020 Research and Innovation Programme under the HORIZON-MSCA-2023-SE-01 JENNIFER3 grant agreement no.101183137, by the Junta de Andalucia (grant No. FQM160) and by University of Tokyo ICRR’s Inter-University Research Program FY2024 (Ref. 2024i-J-001)
Appendix A Comparison to SuSAv2 in bins of $\bf q$ and $Q^{2}$
Figure 21: Comparison to ED-RMF and SuSAv2 for fixed bins in q: Our extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ for ${\rm{}^{12}C}$ for $\bf q$ values of 0.10, 0.148, 0.167, 0.205 and 0.240 GeV versus $\nu$. In all the plots the universal fit for the total (from all processes) ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ is the solid black line and the QE component (including Transverse Enhancement) of the universal fit is the dotted line. The predictions of SuSAv2 Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) are the dashed pink line for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of the ED-RMF Franco-Munoz et al. (2022, 2023) (QE-1p1h including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC model prediction for the 2p2h contribution (solid green line).
Figure 22: Comparison to ED-RMF and SuSAv2 for fixed bins in q: Same as Fig. 21 for ${\bf q}$ values of 0.30, 0.38, 0.475, 0.57 and 0.649 GeV versus $\nu$. The predictions of SuSAv2 Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) are the dashed pink line for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of ED-RMF Franco-Munoz et al. (2022, 2023) (QE-1p1h including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC model prediction for the 2p2h contribution (solid green line).
Figure 23: Comparison to ED-RMF and SuSAv2 for fixed bins in q: Same as Fig. 21 for $\bf q$ values of 0.756, 0.991, 1.619 and 1.921 GeV versus $\nu$. The predictions of SuSAv2 Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) are the dashed pink line for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of ED-RMF Franco-Munoz et al. (2022, 2023) (QE-1p1h including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC model prediction for the 2p2h contribution (solid green line).
Figure 24: Comparison to ED-RMF and SuSAv2 for fixed bins in q: Same as Fig. 21 for $\bf q$ values of 2.4, 2.783, and 3.5 GeV versus $\nu$. The predictions of SuSAv2 Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) are the dashed pink line for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of ED-RMF Franco-Munoz et al. (2022, 2023) (QE-1p1h including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC model prediction for the 2p2h contribution (solid green line).
Figure 25: Comparison to ED-RMF and SuSAv2 for fixed $Q^{2}$ bins: Same as Fig. 21 for $Q^{2}$ values of 0.01 0.02, 0.026, 0.04 and 0.056 GeV2 versus $\nu$. The predictions of SuSAv2 Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) are the dashed pink line for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of ED-RMF Franco-Munoz et al. (2022, 2023) (QE-1p1h including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC model prediction for the 2p2h contribution (solid green line).
Figure 26: Comparison to ED-RMF and SuSAv2 for fixed $Q^{2}$ bins: Same as Fig. 25 for $Q^{2}$ values of 0.093 0.12, 0.16, 0.265 and 0.38 GeV2 versus $\nu$. The predictions of SuSAv2 Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) are the dashed pink line for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of ED-RMF Franco-Munoz et al. (2022, 2023) (QE-1p1h including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC model prediction for the 2p2h contribution (solid green line).
Figure 27: Comparison to ED-RMF and SuSAv2 for fixed $Q^{2}$ bins: Same as Fig. 25 for $Q^{2}$ values of 0.50, 0.8, 1.25, 1.75 and 2.25 GeV2 versus $\nu$. The predictions of SuSAv2 Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) are the dashed pink line for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of ED-RMF Franco-Munoz et al. (2022, 2023) (QE-1p1h including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC model prediction for the 2p2h contribution (solid green line).
Figure 28: Comparison to ED-RMF and SuSAv2 for fixed $Q^{2}$ bins: Same as Fig. 25 for $Q^{2}$ values of 2.75, 3.25 and 3.75 GeV2 versus $\nu$. The predictions of SuSAv2 Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023) are the dashed pink line for QE-1p1h, and solid pink lines for the sum of QE-1p1h and MEC-2p2h processes. The blue lines are the predictions of ED-RMF Franco-Munoz et al. (2022, 2023) (QE-1p1h including nuclear excitations). Since ED-RMF does not account for 2p2h final states, we also show the sum of the ED-RMF prediction for QE-1p1h and the SuSAv2 MEC model prediction for the 2p2h contribution (solid green line).

### v3

@@
 Recent theoretical  Mihaila and Heisenberg (2000); Lovato et al.
-(2016); Cloët et al.
+(2016, 2023a); Cloët et al.
 (2016); Franco-Munoz et al.
@@
 (2020); Carlson et al.
-(2015, 2002) of ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ can be tested by comparing the predictions to experimental data.
+(2015, 2002) calculations of ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ can be tested by comparing the predictions to experimental data.
 We perform extraction of both response functions as functions of (${\bf q},\nu$), as well as ($Q^{2},\nu)$ from an analysis of all available inclusive cross section data for $\rm{}^{12}C$, and compare them to theoretical predictions.
@@
 In this paper we focus on the nuclear excitation, quasielastic (QE) and two nucleon final state (2p2h) regions and compare the extracted $\rm{}^{12}C$ ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ to nuclear theory predictions of the “Green’s Function Monte Carlo” (GFMC) Lovato et al.
-(2016), the “Energy Dependent-Relativistic Mean Field” (ED-RMF) Franco-Munoz et al.
+(2016, 2023a), the “Energy Dependent-Relativistic Mean Field” (ED-RMF) Franco-Munoz et al.
 (2022, 2023), the “Short Time Approximation Quantum Monte Carlo” (STA-QMC) Andreoli et al.
+(2024); Pastore et al.
 (2020), the “Correlated Fermi Gas” (CFG) Bhattacharya et al.
@@
 Recently, ED-RMF has been implementedMcKean (2024) in an update of the neut neutrino MC generator.
-An earlier version of SuSAv2 (QE+2p2h) has been implemented in genie, SuSAv2-inelastic is being implemented in genie, and STA-QMC for $\rm{}^{4}He$ only has been implemented in genie.
+The improved SuSAv2 predictions for the 1p1h and 2p2h channels has been implemented in genie.
+SuSAv2-inelastic is being implemented in genie, and STA-QMC for $\rm{}^{4}He$ has been implemented in genie.
 In this paper we focus on testing theoretical predictions.
@@
 We note that a significant number of cross section measurements were not available in tabular form and required digitization of the cross sections from figures in the publications.
-As summarized in Table 1, the response functions are extracted for a large range of energy transfer $\nu$ at 18 fixed values of $\bf q$ in the range $3.750.1<{\bf q}<$ 3.75  GeV and at 18 fixed values of $Q^{2}$ in the range $3.50<Q^{2}<$ 3.5 GeV2.
+As summarized in Table 1, the response functions are extracted for a large range of energy transfer $\nu$ at 18 fixed values of $\bf q$ in the range $0.1<{\bf q}<3.75$  GeV and at 18 fixed values of $Q^{2}$ in the range $0<Q^{2}<3.5$ GeV2.
 The range of $\nu$ spans the nuclear excitation, QE, resonance and inelastic continuum.
@@
 We also include high precision cross section measurements from Jefferson Lab Hall C experiment E-04-001 Alsalmi et al.
-(2024); Albayrak et al.
 (2025); Alsalmi (2019).
@@
 In the nuclear excitation region (excitation energy less than 50 MeV) we extract ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from fits to $\rm{}^{12}C$} nuclear excitation form factors by Bodek and Christy Bodek and Christy (2023).
-In addition, for $4016<E_{x}<$ 40 MeV, values of ${\cal R}_{L}({\bf q},E_{x})$ and ${\cal R}_{T}({\bf q},E_{x})$ are available from the analysis of Yamaguchi:71 Yamaguchi et al.
+In addition, for $16<E_{x}<40$ MeV, values of ${\cal R}_{L}({\bf q},E_{x})$ and ${\cal R}_{T}({\bf q},E_{x})$ are available from the analysis of Yamaguchi:71 Yamaguchi et al.
 (1971) for five values of $\bf q$ (0.148, 0.167, 0.205, 0.240 and 0.307 GeV).
@@
 I.1 Overview of this analysis
-The Christy-Bodek universal fit to all available cross section measurements for ${\rm{}^{12}C}$ (and Hydrogen and Deuterium) is described in two recent publications Bodek and Christy (2022, 2023) (similar fits were also done for other nuclei).
+The Christy-Bodek 2024 universal fit to all available cross section measurements for $\rm{}^{12}C$} (and Hydrogen and Deuterium) is described in two recent publications Bodek and Christy (2022, 2023) (similar fits were also done for other nuclei).
 The fit includes nuclear elastic form factors, nuclear-excitations, quasielastic scattering, resonance production and inelastic continuum.
@@
 (2014) with Rosenfelder Rosenfelder (1980) Pauli suppression.
-For the QE region the fit also includes parametrizations of a $multiplicative$ “Longitudinal Quenching Factor” at low $\bf q$ and an $additive$ “Transverse Enhancement” contribution, which is significant at intermediate $\bf q$.
+For the QE region the fit also includes parametrizations of a multiplicativemultiplicative “Longitudinal Quenching Factor” at low $\bf q$ and an additiveadditive “Transverse Enhancement” contribution, which is significant at intermediate $\bf q$.
 In the inelastic resonance and continuum region the fit is based on a Gaussian Fermi motion smeared free proton and neutron cross sections with a multiplicative medium modification factor.
@@
   2. 2.
-    Coulomb corrections Gueye et al. (1999) using the Effective Momentum Approximation (EMA) in modeling scattering from nuclear targets. We use $V_{eff}$= 3.10$\pm$ 0.25 MeV for $\rm{}^{12}C$
+[…]
   3. 3.
@@
   9. 9.
-    A $\bf q$ dependent $E_{\mathrm{shift}}^{QE}(\bf q)$ parameter for the QE process to account for the optical potential Bodek and Cai (2019, 2020) of final state nucleons.
+    A $\bf q$ dependent $E_{\mathrm{shift}}^{QE}(\bf q)$ parameter for the QE process to account for the optical potential Bodek and Cai (2019) of final state nucleons.
 The separation energies of a proton and a neutron from $\rm{}^{12}C$ are 16 MeV and 18.7 MeV, respectively.
@@
   14. 14.
+[…]
+  15. 15.
+    There is an apparent shift of the peak of the $\Delta$(1232) to smaller values of $\nu$ which in the 2024 version of the fit is accounted for by an “effective optical potential”Bodek and Cai (2020).
+  16. 16.
     We include QE data at all values of $Q^{2}$ down to $Q^{2}=0.01$ GeV2 ($\bf q$=0.1 GeV) (which were not included in the Bosted-Mamyan fit Bosted and Mamyan (2012).)
-  15. 15.
+  17. 17.
     The relative normalizations between different experiments are extracted from the fit.
-  16. 16.
+  18. 18.
     Data sets which are inconsistent Zeller (1973b); Heimlich et al. (1974) with the world’s data are identified and not included in the analysis (see Table 2).
@@
 Therefore, the fit can also be used to validate nuclear models and tune MC generators for electron and neutrino scattering experiments over a large kinematic range.
-The use of the Christy-Bodek universal fit Bodek and Christy (2022, 2023) in identifying the few data sets that are inconsistent with all the other measurements is illustrated in the top six sub-figures of Fig.
+The use of the Christy-Bodek 2024 universal fit Bodek and Christy (2022, 2023) in identifying the few data sets that are inconsistent with all the other measurements is illustrated in the top six sub-figures of Fig.
 3.
@@
 peaks) in the cross section remain fixed independent of the angle.
-For this reason we perform the analysis in bins of both the excitation energy and bins of the square of the invariant final state mass ($W^{2}=M^{2}+2M\nu-Q^{2}$).
+For this reason we perform the analysis in bins of both the excitation energy and bins of the square of the final state mass ($W^{2}=M^{2}+2M\nu-Q^{2}$).
 We use the results of the analysis in bins of excitation energy for $E_{x}$ less than 50 MeV (a region dominated by nuclear excitations) and the results of the analysis in bins of $W^{2}$ for $E_{x}$ greater than 50 MeV (a region dominated by QE scattering and pion production).
 Afterwards, we convert the $E_{x}$ and $W^{2}$ values in the center of each bin to the corresponding values of $\nu$ in the center of the bin.
-The longitudinal and transverse response functions (defined in section II.3) in the nuclear excitation region are well described by the universal fit Bodek and Christy (2023) to measured transverse and longitudinal form factors for each nuclear excitation.
-The ${\rm{}^{12}C}$ ${\cal R}_{L}/Z^{2}$ and ${\cal R}_{T}/Z^{2}$ measurements published in Yamaguchi:71 Yamaguchi et al.
-(1971) in the nuclear excitation region above the proton separation energy (16$<E_{x}<40$ MeV) are compared to the ${\cal R}_{L}/Z^{2}$ and ${\cal R}_{T}/Z^{2}$ predictions from the fit in Figure 1.
+[…]
 Rosenbluth extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ using data spanning a range of angles from different experiments are valid in the QE and pion production regions.
@@
 For $E_{x}<$16 MeV we use ${\cal R}_{L}$ and ${\cal R}_{T}$ from overall fits Bodek and Christy (2023) to the nuclear excitation form factors (shown as the red solid line in Figure 1).
-For $4016<E_{x}<$ 40 MeV (when available) we use the precise (±$\pm 3\%$) Yamaguchi:71 measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$.
+For $16<E_{x}<40$ MeV (when available) we use the precise ($\pm 3\%$) Yamaguchi:71 measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$.
 In addition, for ${\cal R}_{T}$, we also use electron scattering data at 180 Goldemberg and Barber (1964); Deforest et al.
@@
 0($\gamma^{12}C$) 0 0
-0.010 0.004 0.015 0.100 0.0016$\pm$0.0004 0.063 0.124
-0.020 0.015 0.025 0.148 0.0021$\pm$0.0007 0.124 0.158
-0.026 0.025 0.035 0.167 0.0028$\pm$0.0006 0.158 0.186
-0.040 0.035 0.045 0.205 0.0071$\pm$0.0007 0.186 0.223
-0.056 0.045 0.070 0.240 0.0134$\pm$0.0009 0.223 0.270
-0.093 0.070 0.100 0.300 0.0270$\pm$0.0006 0.270 0.340
-0.120 0.100 0.145 0.380 0.0324$\pm$0.0005 0.340 0.428
-0.160 0.145 0.209 0.475 0.0276$\pm$0.0005 0.428 0.523
-0.265 0.206 0.323 0.570 0.0262$\pm$0.0008 0.523 0.609
-0.380 0.322 0.438 0.649 0.0290$\pm$0.0006 0.609 0.702
-0.500 0.438 0.650 0.756 0.0299$\pm$0.0010 0.702 0.878
-0.800 0.650 1.050 0.991 0.0371$\pm$0.0003 0.878 1.302
-1.250 1.050 1.500 1.619 0.0414$\pm$0.0010 1.302 1.770
-1.750 1.500 2.000 1.921 0.0479$\pm$0.0010 1.770 2.067
-2.250 2.000 2.500 2.213 0.0542$\pm$0.0020 2.067 2.357
-2.750 2.500 3.000 2.500 0.0603$\pm$0.0030 2.357 2.642
-3.250 3.000 3.500 2.783 0.0664$\pm$0.0030 2.642 2.923
-3.750 3.500 4.000 3.500 0.0817$\pm$0.0030 2.923 4.500
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 1: A summary of the 18 bins in $Q^{2}$ (in GeV2) and the 18 bins in $\bf q$ (in GeV). The ${\cal R}_{T}$ (in MeV-1) values for $\nu=\bf q$ are extracted from photo-absorption cross sections.
 Figure 4: Sample ${\cal R}_{L}$ ${\cal R}_{T}$ Rosenbluth plots (before bin centering corrections).
-Figure 5: Nuclear excitation region: Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region ($4016<E_{x}<$ 40 MeV) for ${\rm{}^{12}C}$.
+Figure 5: Nuclear excitation region: Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region ($16<E_{x}<40$ MeV) for $\rm{}^{12}C$}.
 The measurements are from Yamaguchi:71 Yamaguchi et al.
@@
 3 Sealock:89 Sealock et al. (1989) 1.06 0.1000
-Baran:88 Baran et al. (1988) 0.9924 0.0046
+4 Baran:88 Baran et al. (1988) 0.9924 0.0046
 5 Bagdasaryan:88 Bagdasaryan et al. (1988) 0.9878 0.0083
 6 Dai:19 Murphy et al. (2019) 1.0108 0.0053
-7 Arrington9:96 Arrington et al. (1996) 0.9743 0.0133
+7 Arrington:96Arrington et al. (1996) 0.9743 0.0133
 8 Day93 Day et al. (1993) 1.0071 0.0033
-9 Arrington9:99 Arrington et al. (1999) 0.9888 0.0034
+9 Arrington:98 Arrington et al. (1999) 0.9888 0.0034
 10 Gaskell:21 Arrington et al. (2021); Seely et al. (2009) 0.9934 0.0051
 11 Whitney:74 Whitney et al. (1974); Moniz et al. (1971) 1.0149 0.0153
-12 E04-001-2005:24 Alsalmi et al. (2024); Albayrak et al. (2024); Alsalmi (2019); Alsalmi et al. (2025) 0.9981 0.0067
-13 E04-001-2007:24 Alsalmi et al. (2024); Albayrak et al. (2024); Alsalmi (2019); Alsalmi et al. (2025) 1.0029 0.0070
+12 E04-001-2005:24 Alsalmi (2019); Alsalmi et al. (2025) 0.9981 0.0067
+13 E04-001-2007:24 Alsalmi (2019); Alsalmi et al. (2025) 1.0029 0.0070
 14 Gomez:74 Gomez et al. (1994); Res 1.0125 0.0149
@@
 22 Goldemberg64 Goldemberg and Barber (1964)(180) 1.100 0.1000
-23 Defores:t65 Deforest et al. (1965)(180) 0.9 0.1000
+23 Deforest:65 Deforest et al. (1965)(180) 0.9 0.1000
 24 Mainz:2024Mihovilovič et al. (2024) (not used) 1.03 0.02
-25 CLAS-e4nu:2023Isaacson et al. (2023) (not used) 0.85 0.02
+25 CLAS-e4nu:2023Isaacson et al. (2023) 0.85 0.02
+26 Garino::1992Garino et al. (1992) (not used) 1.00 0.02
+27 Ricco:1968:Ricco:1968bsq (not used) - -
 30 Donnelly:68 Donnelly et al. (1968); Donnelly (1970) (not used) - -
 31 Zeller:73 Zeller (1973b); Heimlich et al. (1974) (not used) - -
-Table 2: A summary table of the $\rm{}^{12}C$ data sets used in the Christy-Bodek universal fit and in this analysis.
+Ahrens:75 Ahrens et al. (1975)(photoabsorption) - -
+(not-included-in-fit)
+Carrasco:89 Carrasco and Oset (1992)(photoabsorption) - -
+Bianchi:95 Bianchi et al. (1996)(photoabsorption) - -
+Bezic:69 Bezić et al. (1969)(photoabsorption) - -
+(not-included-in-fit)
+Table 2: A summary table of the $\rm{}^{12}C$ data sets used in the Christy-Bodek 2024 universal fit and in this analysis.
 Shown are the normalization factors and uncertainties.
@@
 For scattering from a nuclear target such as carbon, the excitation energy $E_{x}$ is given by $E_{x}=\nu-\nu_{elastic}$ where
-$\nu_{elastic}=E_{0}-\frac{E_{0}}{1+2E_{0}{\sin}^{2}\frac{\theta}{2}/M_{A}},$ (4)
-and $M_{A}$ is the mass of the nuclear target. For carbon $M_{A}=12u=11.178$ GeV (1 u = 931.502 MeV). Or equivalently
-$E_{x}=\nu-\frac{Q^{2}_{elastic}}{2M_{A}},$ (5)
-where $Q^{2}_{elastic}$ is the $Q^{2}$ for elastic scattering from a carbon nucleus for incident energy $E_{0}$ and scattering angle $\theta$.
+$\nu_{elastic}=E_{0}-\frac{E_{0}}{1+2E_{0}{\sin}^{2}\frac{\theta}{2}/M_{A}}=\frac{Q^{2}_{elastic}}{2M_{A}},$ (4)
+where $Q^{2}_{elastic}$ is the $Q^{2}$ for elastic scattering from a carbon nucleus for incident energy $E_{0}$ and scattering angle $\theta$ and $M_{A}$ is the mass of the nuclear target.
+For carbon $M_{A}=12u=11.178$ GeV (1 u = 931.502 MeV).
+The excitation energy is
+$E_{x}=\nu-\nu_{elastic}$ (5)
+Alternatively, one can describe nuclear excitations in terms of the mass of the excited carbon nucleus.
+$W^{2}_{A}=M^{2}_{A}+2M_{A}\nu-Q^{2},$ (6)
 II.1 Description in terms of longitudinal and transverse virtual photon cross sections
@@
 Bottom: The response function ${\cal R}_{T}({Q^{2}=0},\nu)$ extracted from photo-absorption cross sections using eq.
-24.
+25.
 II.3.1 Extraction of response functions from photo-absorption cross sections
@@
 (1996); Muccifora et al.
-(1999) on 12C is shown in the top panel of Fig 6, while ${\cal R}_{T}(Q^{2}=0,\nu)$ extracted via Eq 24 is shown in the bottom panel.
-In electron scattering $Q^{2}$=${\bf q}^{2}+\nu^{2}$ must be $\geq 0$.
-Therefore, for a fixed $\bf q$ the maximum $\nu$ ($\nu^{max}$) occurs when $\nu=\bf q$.
+(1999) on 12C is shown on the top panel of Fig 6.
+The dashed red line is the expectation from the Quasi-deuteron model Plujko et al.
+(2018)(scattering from short-correlation neutron-proton pairs).
+The Values of ${\cal R}_{T}(Q^{2}=0,\nu)$ extracted via Eq 25 is shown in the bottom panel.
+In electron scattering $Q^{2}$${\bf q}^{2}+\nu^{2}$ must be $\geq 0$.
+Therefore, for a fixed $\bf q$ the maximum $\nu$ ($\nu^{max}$) occurs when $\nu=\bf q$ and
 ${\cal R}_{T}({\bf q,}\nu=\bf q)$ $=$ $\frac{\bf q}{2\pi^{2}\alpha}\sigma_{\gamma}(\bf q).$ (27)
 In photo-absorption the cross section is dominated by the Giant Dipole Resonance, quasi-deuterons Tavares and Terranova (1992) and pion production processes (the cross section for the quasielastic process is zero).
-II.4 Experimental extraction of response functions
 We use a form of the Rosenbluth method introduced by Jourdan Jourdan (1996, 1995) to separate ${\cal R}_{L}({Q^{2}},\nu)$ and ${\cal R}_{T}({Q^{2}},\nu)$ as follows:
@@
 However, when Coulomb corrections are included, the above expressions are modified and all the parameters are replaced with “effective parameters” as described below.
-II.5 Electron scattering at 180
+II.4 Electron scattering at 180
 Including Coulomb corrections the response function ${\cal R}_{T}({Q_{eff}^{2}},\nu)$ can be extracted directly from the electron scattering cross section at 180 using the following expression:
 ${\cal R}_{T}({Q_{eff}^{2}},\nu)$ $=$ $\left(\frac{E_{0}}{E_{0,eff}}\right)^{2}\frac{{Q_{eff}}^{4}}{4\alpha^{2}E_{eff}^{{}^{\prime}2}}\frac{d\sigma}{d\nu d\Omega},$ (31)
-where the $eff$ subscript denotes Coulomb corrected quantities described below.
-II.6 ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region
-As mentioned earlier, in the nuclear excitation region $4016<E_{x}<$ 40 MeV we use the precise (±$\pm 3\%$) Yamaguchi measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$ (where available).
+where the effeff subscript denotes Coulomb corrected quantities described below.
+II.5 ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region
+As mentioned earlier, in the nuclear excitation region $16<E_{x}<40$ MeV we use the precise ($\pm 3\%$) Yamaguchi measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$ (where available).
 For $E_{x}<16$ MeV we extract ${\cal R}_{L}$ and ${\cal R}_{T}$ from fits Bodek and Christy (2023) to the nuclear excitation form factors, and performing a Gaussian smearing with the width determined by the experimental resolution.
@@
 The experimentally determined $V_{eff}$ for carbon is $V_{eff}$= 3.1 $\pm$ 0.25 MeV.
-Including Coulomb corrections we define:
+Including Coulomb corrections (CC) we define:
 $E_{0,eff}$ $=$ $E_{0}+V_{eff}$ (35)
@@
 $W^{2}_{eff}$ $=$ $M^{2}+2M\nu-Q^{2}_{eff}$
+$E_{x}^{cc}$ $=$ $E_{x}$
+$W^{2}_{A-eff}$ $=$ $M^{2}_{A}+2M_{A}\nu-Q^{2}_{eff}$
+$\epsilon^{CC}$ $=$ $\left[1+2\left(1+\frac{\nu^{2}}{Q_{eff}^{2}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
 The response functions are calculated with $Q^{2}=Q^{2}_{eff}$ and $E^{\prime}=E^{\prime}_{eff}$.
@@
 Therefore
-$\epsilon^{CC}$ $=$ $\left[1+2\left(1+\frac{\nu^{2}}{Q^{2}_{eff}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
 $\Sigma^{\prime}({Q^{2}_{eff}},\nu)=H^{CC}\frac{d\sigma}{d\nu d\Omega}$
@@
 IV Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$
-From the Christy-Bodek overall fit we extract the relative normalization ($N_{i}$) of the various data sets as given in Table 2.
-For low excitation energy ($E_{x}<$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of excitation energy as described below.
-Similarly for ($E_{x}>$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of the square of the final state invariant mass $W^{2}$.
-We use the fit to apply $Q^{2}$ bin centering corrections ($C_{i}$) to all extracted value off $\Sigma_{i}^{\prime}$ that are within the $Q_{eff}^{2}$ range for each $Q_{eff}^{2}$ and of $W_{eff}^{2}$ bin such that they reflect the values at $Q_{enter}^{2}$, $W_{center}^{2}$, and $\epsilon_{center}$ for each cross section measurement,
-$C_{i}$ $=$ $\frac{\epsilon{\cal R}^{fit}_{L}({Q_{center}^{2})}+\frac{1}{2}(\frac{{\bf q_{center}}}{Q_{center}})^{2}{\cal R}^{fit}_{T}({Q_{center}^{2}})}{\epsilon{\cal R}^{fit}_{L}({Q_{eff}^{2})}+\frac{1}{2}(\frac{{\bf q_{eff}}}{Q_{eff}})^{2}{\cal R}^{fit}_{T}({Q_{eff}^{2}})}.$
+From the Christy-Bodek 2024 overall fit we extract the relative normalization ($N_{i}$) of the various data sets as given in Table 2.
+For low excitation energy ($E_{x}<$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of the final state excited nuclear mass $W^{2}_{A}$, as described below.
+Similarly for ($E_{x}>$ 50 MeV) bin centering corrections are minimized if we extract the response functions as functions of the square of the final state mass defined in terms of the mass of the nucleon $W^{2}$.
+We use the fit to apply $Q^{2}$ bin centering corrections ($C_{i}$) to all extracted value of $\Sigma_{i}^{\prime}$ that are within the $Q_{eff}^{2}$ range for each $Q_{eff}^{2}$ and of $W_{eff}^{2}$ bin such that they reflect the values at $Q_{enter}^{2}$, $W_{center}^{2}$, and $\epsilon_{center}$ for each cross section measurement,
+IV.1 Bin centering corrections for fix $Q^{2}$ bins
+$C_{i}=\frac{\epsilon_{center}{\cal R}^{fit}_{L}+\frac{1}{2}(\frac{{\bf q_{center}}}{Q_{center}})^{2}{\cal R}^{fit}_{T}}{\epsilon^{cc}{\cal R}^{fit}_{L}(W^{2}_{eff},{Q_{eff}^{2})}+\frac{1}{2}(\frac{{\bf q_{eff}}}{Q_{eff}})^{2}{\cal R}^{fit}_{T}({W^{2}_{eff},Q_{eff}^{2}})},$
+For excitation energies above 50 MeV we use $W^{2}$ based bin centering corrections.
+${\cal R}^{fit}_{L}={\cal R}^{fit}_{L}(W^{2}_{center},{Q_{center}^{2})}$
+${\cal R}^{fit}_{T}={\cal R}^{fit}_{T}(W^{2}_{center},{Q_{center}^{2})}$
+$\nu_{center}=\frac{W^{2}_{center}-M^{2}+Q^{2}_{center}}{2M}$ (37)
+and
+$\epsilon_{center}=\left[1+2\left(1+\frac{\nu^{2}_{center}}{Q_{center}^{2}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
 After the application of the bin centering corrections to all extracted values $\Sigma^{\prime}_{i}$ in each $Q_{eff}^{2}$ bin we can assume that extracted values of $\Sigma_{i}$ are at $Q^{2}_{center}$.
-When we apply these corrections, we keep the final state invariant mass $W^{2}$ fixed for $\nu>$ 50 MeV, and only correct for the change in $Q_{eff}^{2}$.
-For $\nu<$ 50 MeV we use the excitation energy $E_{x}$ instead of $W^{2}$.
-For $E_{x}$ bins at fixed values of $Q^{2}$ the value of $\nu$ in the bin center is extracted from value of $E_{x}$ in the bin center using the following expression.
-$\nu$ $=$ $E_{x}+\frac{Q^{2}}{2M_{A}},$ (36)
-where $M_{A}$ is the mass of the nuclear target (for carbon $M_{A}$=11.178 GeV).
-For $W^{2}$ bins at fixed values of $Q^{2}$, the value of $\nu$ at the bin center is extracted from value of $W^{2}$ in the bin center using the following expression:
-$W^{2}$ $=$ $M^{2}+2M\nu-Q^{2},$ (37)
-For $E_{x}$ bins at fixed values of $\bf q$ the value of $\nu$ at the bin center is extracted from value of $E_{x}$ in the bin center by solving the following quadratic equation:
-$\nu$ $=$ $E_{x}+({\bf q}^{2}+\nu^{2})/(2M_{A})$ (38)
-For $W^{2}$ bins at fixed values of $\bf q$ the value of $\nu$ in the bin center is extracted from value of $W^{2}$ in the bin center by solving the following quadratic equation:
-$W^{2}$ $=$ $M^{2}+2M\nu-({\bf q}^{2}+\nu^{2}).$ (39)
-IV.1 Comparison to Previous Measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
+When we apply these corrections, we keep the final state mass $W^{2}$ fixed for $\nu>$ 50 MeV, and only correct for the change resulting from changing $Q_{eff}^{2}$ to $Q^{2}_{center}$, which also results in changing $\nu$ to $\nu_{center}$ and $\epsilon_{cc}$ to $\epsilon_{center}$.
+For $Ex<$ 50 MeV at fixed values of $Q^{2}$ the value of $\nu_{center}$ is
+$\nu_{center}=E_{x}+\frac{Q^{2}_{center}}{2M_{A}}$ (39)
+where $M_{A}$ is the mass of the nuclear target (11.178 GeV for carbon). and
+$\epsilon_{center}=\left[1+2\left(1+\frac{\nu^{2}_{center}}{Q_{center}^{2}}\right){\tan}^{2}\frac{\theta}{2}\right]^{-1}$
+$W^{2}_{center}=M^{2}+2M\nu_{center}-Q^{2}_{center}$
+${\cal R}^{fit}_{L}={\cal R}^{fit}_{L}(W^{2}_{center},{Q_{center}^{2})}$
+${\cal R}^{fit}_{T}={\cal R}^{fit}_{T}(W^{2}_{center},{Q_{center}^{2})}.$ (40)
+For fixed $Q^{2}$ bins, after bin centering, all of the Rosenbluth fits are done as a function of of $\epsilon_{center}$, and the extracted ${\cal R}_{L}$ and ${\cal R}_{T}$ are plotted as a function of $\nu_{center}$.
+IV.2 Bin centering corrections for fixed $\bf q$ bins
+For $E_{x}>$ 50 MeV. we solve the following quadratic equation for $\nu_{center}$.
+$W^{2}_{center}$ $=$ $M^{2}+2M\nu_{center}-({\bf q}^{2}_{center}-\nu^{2}_{center}),$
+which yields
+$Q^{2}_{center}={\bf q}^{2}_{center}-\nu^{2}_{center,}$
+and then follow up with equations IV.1 and IV.1.
+For $Ex<$ 50 MeV at fixed $\bf q$ beins We solve the following quadratic equation for $\nu_{center}$.
+$\nu_{center}$ $=$ $E_{x}+\frac{{\bf q}^{2}-\nu^{2}_{center}}{2M_{A}}$ (41)
+which yields
+$Q^{2}_{center}={\bf q}^{2}_{center}-\nu^{2}_{center},$
+$W^{2}_{center}=M^{2}+2M\nu_{center}-Q^{2}_{center},$
+and then follow up with equation IV.1.
+IV.3 Comparison to Previous Measurements of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
 A comparison between our extraction of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ (red diamonds) to previous extractions are shown in Fig.
@@
 A later extraction by Jourdan:96 was also performed for three slightly different values of momentum transfer (${\bf q}=0.30$, 0.38 and 0.57 GeV).
-In the Jourdan analysis the same Barreau:82 cross sections used but with additional cross section data from SLAC.
+In the Jourdan analysis the same Barreau:82 cross sections used, with additional cross section data from SLAC.
 Our extraction is more extensive because our data sample is much larger than what was available in 1996 (as summarized in Table 2).
@@
 Note that the more recent Kharkov analysis has larger errors and was performed only for ${\bf q}=0.3$ GeV.
-There is good agreement between our extraction at $Q^{2}$=0.093 GeV2222 to the extraction by Baran:1988 Baran et al.
+There is good agreement between our extraction at $Q^{2}$=0.093 GeV22 and the extraction by Baran:1988 Baran et al.
 (1988) at $Q^{2}$=0.1 GeV2 (light green circles) in the region of the $\Delta$(1238) resonance as seen in Figure 7(B).
+IV.4 Systematic Errors
+We estimate an overall systematic error of 4% in the extracted values of the longitudinal and transverse response functions.
+This error is anti-correlated between ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ .
 V Comparison to theoretical predictions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$
@@
 The values of ${\cal R}_{T}(\bf q,\nu)$ at $\nu$ = $\bf q$ ($Q^{2}$ = 0) are extracted from photo-absorption measurements.
-${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ extracted from the Christy-Bodek universal fit (which includes nuclear excitations) are shown as solid black lines, and the QE contributions to the fit (including TE) are the dashed black lines.
+${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ extracted from the Christy-Bodek 2024 universal fit (which includes nuclear excitations) are shown as solid black lines, and the QE contributions to the fit (including TE) are the dashed black lines.
 Figure 8: Comparison to ED-RMF: Our extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ for $\rm{}^{12}C$} for $\bf q$ values of 0.10, 0.148, 0.167, 0.205, and 0.240 GeV versus $\nu$.
@@
     “Green’s Function Monte Carlo” (GFMC)Lovato et al.
-(2016) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states.
+(2016, 2023a) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states.
 GFMC predictions are only available for $0.3\leq{\bf q}\leq 0.57$ GeV.
@@
 (2024) which models the contribution of 1-body and two body-currents to the response functions for single nucleon (1p1h) and two nucleon (2p2h) final states (for both electron and neutrino scattering processes).
-Currently, the STA-QMC predictions are only valid for $0.3\leq{\bf q}\leq 0.76$ GeV.
+Currently, the STA-QMC predictions are valid for $0.3\leq{\bf q}\leq 0.65$ GeV.
   5. 5.
@@
 In addition, the SuSAv2-inelastic model is currently being implemented in genie for electrons and neutrinos.
-The SuSaA2 2p2h channel has been implemented (only for neutrinos) in NuWro,  andSTA-QMC been implementedBarrow et al.
-(2021) in genie  but only for $\rm He^{4}$ (though an hadron tensor interfaceGardiner (2019)), and the achilles generator is being developed for use in future neutrino experiments.
+The SuSAv2 2p2h channel has been implemented (only for neutrinos) in NuWro,  the achilles generator is being developed for use in future neutrino experiments, and STA-QMC been implementedBarrow et al.
+(2021) in genie  for $\rm He^{4}$ (though an hadron tensor interfaceGardiner (2019)).
 V.1 Comparison to the predictions of GFMC, STA-QMC, and ED-RMF
@@
 The three theoretical calculations include contributions from both 1-body and 2-body (1b+2b) currents.
-In these three calculations there is enhancement of the calculated transverse response function ${\cal R}_{T}(\bf q,\nu)$ The transverse enhancement (TE) primarily originates from the interference Simons et al.
+In these three calculations there is enhancement of the calculated transverse response function ${\cal R}_{T}(\bf q,\nu)$ .
+The transverse enhancement (TE) primarily originates from the interference Carlson et al.
+(2002); Simons et al.
+(2022); Andreoli et al.
 (2024) between 1b and 2b currents which result in 1p1h.
 (1 nucleon 1 hole) final states.
-The contribution of 2b currents to ${\cal R}_{T}(\bf q,\nu)$ which results in 2p2h (2 nucleons 2 holes) final states are not included ED-RMF but are included in GFMC and STA-QMC.
-Note that final states with pions are not included in any of the three calculations (pion production is only included in the improved version of SuSAv20).
+The contribution of 2b currents to ${\cal R}_{T}(\bf q,\nu)$ which results in 2p2h (2 nucleons 2 holes) final states is not included ED-RMF but are included in GFMC and STA-QMC.
+Note that final states with pions is not included in any of the three calculations (pion production is only included in the improved version of SuSAv2).
 The GFMC predictions are shown as the pink lines in Fig.
@@
 Consequently the response functions below $\bf q$=0.3 GeV and above $\bf q$=0.57 are not available for GFMC.
-The predictions for $\rm{}^{12}C$ (shown as the light green lines in Fig.
+In addition, there are large uncertainties in the predictions near threshold (small $\nu$).
+The STA-QMC predictions for $\rm{}^{12}C$ (shown as the light green lines in Fig.
 9-10) are provided for $0.3\leq{\bf q}\leq 0.76$ GeV.
 By its nature STA-QMC only works for $\bf q>$ 0.3 GeV.
-The STA-QMC predictions for for $0.3\leq{\bf q}\leq 0.6$ GeV are in agreement with our measurements.
+The STA-QMC predictions for $0.3\leq{\bf q}\leq 0.6$ GeV are in agreement with our measurements.
 Presently, STA-QMC does not fully account for relativistic effects at the vertex.
-Therefore, although STA-QMC calculations have been provided for $\bf q$=0.76 GeV (shown in Fig.
+Therefore, although STA-QMC calculations have been provided for up to $\bf q$=0.76 GeV (shown in Fig.
 10), it is expected that that they will be less accurate in the high $\bf q$ region where relativistic effects become increasingly more relevant.
-Although the STA-QMC calculations are based on the same many-nucleon Hamiltonian and electromagnetic currents as GFMC, the STA-QMC calculations are in better agreement with our measurements than GFMC.
+The STA-QMC calculations are based on the same many-nucleon Hamiltonian and electromagnetic currents as GFMC.
+Since STA-QMC does not have any knowledge of the correct threshold behavior or low-energy properties of the system, the correct threshold behavior is imposed as a constraintPastore et al.
+(2020).
 The ED-RMF theoretical predictions are shown as the blue lines.
@@
 This is expected because the ED-RMF theoretical calculations for ${\cal R}_{T}(\bf q,\nu)$ do not include 2-body current processes with 2p2h final states or Fermi smeared pion production processes.
-The STA-QMC calculation includes both single and two nucleon final states but is only valid over a more restricted kinematic range and does not account for nuclear excitations.
+The STA-QMC calculation includes both single and two nucleon final states but is valid over a more restricted kinematic range and does not account for nuclear excitations.
 The calculations of both ED-RMF and STA-QMC are directly applicable to the same kinematic regions for neutrino scattering.
@@
 (2019)).
-At low energies of the final nucleon (up to about 20 MeV in the continuum) the ED-RMF potential is the same potential employed to compute the momentum distributions for the initial bound nucleons an at larger energies it approaches the behavior of a phenomenological optical potentials.
+At low energies of the final nucleon (up to about 20 MeV in the continuum) the ED-RMF potential is the same potential employed to compute the momentum distributions for the initial bound nucleons and at larger energies it approaches the behavior of a phenomenological optical potentials.
 This ensures orthogonality between the initial and final states at these low energies and thus the reliability of the cross-section computed in this region.
@@
 The decays of the 12.71 and 16.11 MeV excitations (which are below the proton and neutron separations energy) are primarily via the the emission of $\alpha$ particles, while for the 20.62 MeV excitation the branching ratio to decay via proton emission is a factor of 2 higher than the decay via $\alpha$ emission.
-Note that Christy-Bodek fit to inclusive electron scattering data describes all nuclear excitations irrespective of their decay modes.
+Note that Christy-Bodek 2024 fit to inclusive electron scattering data describes all nuclear excitations irrespective of their decay modes.
 Figure 12: Comparison to CFG, and NuWro for fixed q bins: Our extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ for $\bf q$ values of 0.10, 0.148, 0.167, 0.205, and 0.240 GeV versus $\nu$.
@@
 The two-body currents and interference between one-body and 2-body currents are responsible for the enhancement of the transverse response functions.
-Therefore, in electron scattering mode NuWro is expected to underestimate the transverse response function in the QE region.
-and the comparisons with NuWro are only valid for ${\cal R}_{L}(\bf q,\nu)$
+Therefore, in electron scattering mode NuWro is expected to underestimate the transverse response function and overestimate the longitudinal response function in the QE region
 In addition, in electron scattering mode NuWro does not include pion production processes.
@@
 However, since NuWro-SF-FSI also does not include a model for nuclear excitations, the overestimate of ${\cal R}_{L}(\bf q,\nu)$ partially compensates for the missing nuclear excitations in the longitudinal channel.
-The direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek universal fit can be used to extract corrections (e.g.
+The direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek 2024 universal fit can be used to extract corrections (e.g.
 tuning) that will bring the predictions into better agreement with the data as discussed in section 3.
@@
 As seen in Fig.12 - 15 the model breaks down at low momentum transfers and works somewhat better at higher momentum transfers.
-Here also, the direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek universal fit can be used to extract corrections (e.g.
+Here also, the direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek 2024 universal fit can be used to extract corrections (e.g.
 tuning) that will bring the predictions into better agreement with the data as discussed in section 3.
@@
 For ${\cal R}_{T}(\bf q,\nu)$ this is expected because achilles  also include the contribution of 2-body currents and the interference between 1-body and 2-body currents leading to 1-nucleon final states.
-The predictions for ${\cal R}_{T}(\bf q,\nu)$ are lower than the QE+TE component of Christy-Bodek fit because the transverse enhancement resulting in 2-nucleon final states (2p2h) are not included in the model.
+The predictions for ${\cal R}_{T}(\bf q,\nu)$ are lower than the QE+TE component of Christy-Bodek 2024 fit because the transverse enhancement resulting in 2-nucleon final states (2p2h) are not included in the model.
 Figure 16: Comparison to ED-RMF and NuWro for fixed $Q^{2}$ bins: Same as Fig.
@@
 As expected, the NuWro predictions for ${\cal R}_{T}(\bf q,\nu)$ are underestimated because the electron mode of NuWro does not include the contribution of Meson Exchange Currents (which are included in the neutrino version of NuWro).
-Therefore, for NuWro the relevant validations are only for ${\cal R}_{L}(\bf q,\nu)$
 At low $Q^{2}$ we find that the NuWro-SF-FSI predictions are in better agreement with the data than the predictions of NuWro-SF.
@@
 However, at high values of q and $Q^{2}$ (e.g.
-$\bf q>$1.6 GeV and $Q^{2}>$ 1.25 GeV) the SuSAv2 MEC-2p2h model predictions for the longitudinal response functions have unphysical peaks.
+$\bf q>$1.6 GeV and $Q^{2}>$ 1.25 GeV2) the SuSAv2 MEC-2p2h model predictions for the longitudinal response functions have unphysical peaks.
 This could be because the parameterization of the 2p2h RFG-based nuclear responses from Ruiz-Simo et al.Ruiz Simo et al.
 (2017) included in this implementation of the SuSAv2-MEC model is valid only for ${\bf q}<2$ GeV and $\nu<2$ GeV (although it can be extended to higher kinematics).
-In contrast the SuSAv2-MEC-2p2h implemented in genie  uses interpolation methods and therefore is more accurate than the parameterization formulae used in this implementation.
+Note that the SuSAv2-MEC-2p2h implemented in genie  uses interpolation methods and therefore is more accurate than the parameterization formulae used in this implementation.
 This can introduce important differences at low kinematics ($\bf q$}< 0.3 GeV) and also at very high kinematics ($\bf q$}>1.5 GeV).
+VI.3 Resonance
 Since the ED-RMF-QE-1p1h predictions are in better agreement with our measurements than the predictions of SuSAv2-QE-1p1h, we also show the sum of the predictions of ED-RMF-QE-1p1h and the predictions of SuSAv2-MEC-2p2h as the solid green lines in the plots in Appendix A.
 Combining the ED-RMF-QE-1p1h predictions with the SuSAv2-MEC-2p2h predictions provides a better description of the response function for both single nucleon (from QE and nuclear excitations) and two nucleon final states.
-Model Currents Final State Validity RL RT high $\bf q$ low $\bf q$
-ED-RMF 1b,2b 1p1h, nucl. exci. $all\leavevmode\nobreak\ \bf q$ OK OK 2p2h model 2p2h model
-QMC no 2p2h need 2p2h need 2p2h needed needed
-model model use SuSAV2 ? use SuSAv2?
-SuSAV2 1b 1p1h + $all\leavevmode\nobreak\ \bf q$ low $\nu$ unphysical low $\nu$ unphysical OK need RL quench.
-RMF Scaling for 2p2h model low q high low q low need more RT
-function 1p1h no nucl exc. need quench
-STA-QMC 1b,2b 1p1h, 2p2h $0.650.3<{\bf q}<$ 0.65 OK OK relativistic analytic extrapol.
-no nucl exc. only corr. needed needed
-NuWro SF-FSI 1b 1p1h all q too high too low needs needs
+Model Currents Final State Available for RL RT Large $\bf q$ Small $\bf q$
+ED-RMF 1b,2b 1p1h, nucl. exci. $all\leavevmode\nobreak\ \bf q$ best 1p1h best 1p1h best 1p1h best 1p1h
+QMC 1b+2b best model 2p2h small need 2p2h need 2p2h need 2p2h
+no 2p2h for 1p1h use SuSAv2 use SuSAv2 use SuSAv2 use SuSAv2
+SuSAv2 1p1h 1p1h $all\leavevmode\nobreak\ \bf q$ low $\nu$ unphysical low $\nu$ unphysical OK needs RL
+RMF Scaling only 1b only 1b 1p1h no 1b-2b low q model-high low q model-low quench
+function 2p2h no nucl exc. interference needs 1p1h needs 1p1h needs RT
+1b+2b 2p2h OK 2p2h OK quench or enhancement enhancement
+use ED-RMF? use ED-RMF? use ED-RMF?
+STA-QMC 1b,2b 1p1h + 2p2h. $0.3\leq{\bf q}<\leq.65$ OK OK relativistic analytic extrapol.
+no nucl exc. $0.3\leq{\bf q}<\leq.65$ $0.3\leq{\bf q}\leq 0.65$ corr. needed needed
+NuWro SF-FSI 1b 1p1h all q RL high RT low needs needs
 e-mode spectral no nucl exc. needs needs 2p2h 2p2h
 function no 2p2h quenching enhancement model model
-achilles 1b,2b 1p1h ${\bf q}>$0.5 OK OK RL RT need other
-QMC spectral no 2p2h. only need need need low q models
-function no nucl exc. 2p2h model 2p2h model 2p2h model for 1p1h, 2p2h
-GFMC 1b,2b 1p1h, 2p2h $0.570.3<{\bf q}<$ 0.57 low $\nu$ unphysical OK CPU intensive CPU intensive
-QMC no nucl. exc. only q=0.57 too high not possible not possible
-CFG 1b 1p1h,2p2h all q poor poor poor poor
+achilles 1b,2b 1p1h $\bf q$}>0.5 OK OK needs need other
+QMC spectral no 2p2h. need need 2p2h model low q models
+function no nucl exc. 2p2h model 2p2h model for RL RT for 1p1h, 2p2h
+GFMC 1b,2b 1p1h+ 2p2h $0.3\leq{\bf q}\leq 0.57$ low $\nu$ unphysical OK CPU intensive CPU intensive
+QMC no nucl. exc. q=0.57 high $0.3\leq{\bf q}<\leq.57$ not possible not possible
+CFG 1b 1p1h+2p2h all q poor poor poor poor
 Correlated no nucl. exc. agreement agreement agreement agreement
@@
 Table 3: A summary of comparisons of $\rm{}^{12}C$ ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ to theoretical predictions (q units are in GeV).
-Because our extractions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from all available data electron scattering data span a large range of $\bf q$ and $Q^{2}$, the comparisons can be used extract $\bf q$ dependent corrections to nuclear models and neutrino/electron MC generators to achieve better agreement our measurement s (as has been done in the Christy-Bodek fit).
+Because our extractions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from all available data electron scattering data span a large range of $\bf q$ and $Q^{2}$, the comparisons can be used extract $\bf q$ dependent corrections to nuclear models and neutrino/electron MC generators to achieve better agreement our measurement s (as has been done in the Christy-Bodek 2024 fit).
 For example, in the QE and nuclear excitation region:
@@
 Tuning these parameters in SuSAv2 would bring the position of the QE peaks into better agreement with the measurements at low $\bf q$ and $Q^{2}$.
-In addition, introducing a q-dependence in the RMF-based SuSAv2 scaling functions (which are roughly constant for ${\bf q}>$ 0.3 MeV, but not for smaller values of q) can be used to incorporate into SuSAV2 the the low-energy RMF nuclear effects that break the scaling behavior at low q, which will modify the shape and magnitude of the SuSAv2 QE predictions at that region.
+In addition, introducing a q-dependence in the RMF-based SuSAv2 scaling functions (which are roughly constant for $\bf q$}> 0.3 GeV, but not for smaller values of q) can be used to incorporate into SuSAv2 the the low-energy RMF nuclear effects that break the scaling behavior at low q, which will modify the shape and magnitude of the SuSAv2 QE predictions at that region.
   2. 2.
-    The location of the $\Delta$(1232) is also sensitive to removal energy and and final state optical potential.
-We observe that the shift of the peak of the $\Delta$(1232)  Bodek and Cai (2020) to lower values of $\nu$ is larger than the shift of the QE peak.
-This is unexpected because the optical potential for the $\Delta$(1232) is about 2/3 of the optical potential for the nucleonEricson and Weise (1988).
-In SuSAv2 this larger shift is accounted for by the 2p2h-MEC contribution to the cross section in the dip region.
-In gibuu this shift is accounted for by a nuclear pion production continuumMosel and Gallmeister (2024).
-In the Christy-Bodek fit this additional shift is parametrized in terms of a larger “effective optical potential” Bodek and Cai (2020) for the $\Delta$(1232) and inelastic pion production processes.
-Alternatively, the treatment of the real and imaginary parts of the $\Delta$ propagator in 2p2h and resonance channels and $\Delta$ in-medium modifications can also be tuned or modified to achieve better agreement with electron scattering data.
+    The magnitude of ${\cal R}_{L}(\bf q,\nu)$ can tuned be via a $\bf q$ dependent longitudinal quenching factor, and the amplitude of ${\cal R}_{T}(\bf q,\nu)$ can be tuned via an $\bf q$ dependent additive transverse enhancement contribution (as done in the Christy-Bodek 2024 fitBodek and Christy (2022)).
+Including and tuning such parameters in SuSAv2 would bring the magnitude of ${\cal R}_{L}^{QE}(\bf q,\nu)$ and ${\cal R}_{T}^{QE}(\bf q,\nu)$ predictions into better agreement with the measurements.
   3. 3.
-    The magnitude of ${\cal R}_{L}(\bf q,\nu)$ can tuned be via a $\bf q$ dependent longitudinal quenching factor, and the amplitude of ${\cal R}_{T}(\bf q,\nu)$ can be tuned via an $\bf q$ dependent additive transverse enhancement contribution (as done in the Christy-Bodek fitBodek and Christy (2022)).
-Including and tuning such parameters in SuSAv2 would bring the magnitude of ${\cal R}_{L}^{QE}(\bf q,\nu)$ and ${\cal R}_{T}^{QE}(\bf q,\nu)$ predictions into better agreement with the measurements.
-  4. 4.
     In ED-RMF nuclear excitations (which decay to single nucleons in the final state) occur via the interactions of final state nucleons with the Energy Dependent Mean Field.
 In the other calculations, a model describing nuclear excitations (with excitation energies above the proton separation energy) can be added.
-  5. 5.
-    A model describing nuclear excitations (with excitation energies below the proton separation energy) which decay via $\gamma$ ray emission is included in the inclusive Christy-Bodek fit Bodek and Christy (2023).
+  4. 4.
+    A model describing nuclear excitations (with excitation energies below the proton separation energy) which decay via $\gamma$ ray emission is included in the inclusive Christy-Bodek 2024 fit Bodek and Christy (2023).
 At present none of the theoretical predications include nuclear excitations which decay via $\gamma$ emission.
@@
 Since the extracted response functions cover a large range of $Q^{2}$ and $\nu$, they can be readily used to validate nuclear models as well as validate and tune Monte Carlo generators for electron and neutrino scattering experiments.
-Alternatively, the Christy-Bodek global fit to all electron scattering data on $\rm{}^{12}C$ can also be used to validate nuclear models and also validate and tune Monte Carlo generators for electron and neutrino scattering experiments over a larger region of $\nu$.
+Alternatively, the Christy-Bodek 2024 global fit to all electron scattering data on $\rm{}^{12}C$ can also be used to validate nuclear models and also validate and tune Monte Carlo generators for electron and neutrino scattering experiments over a larger region of $\nu$.
 In this paper we focus on comparison to models in the nuclear excitation region, and for single nucleon (QE-1p1h) and two nucleon (2p2h) final states.
@@
 In our next publication (in which the resonance region will be investigated) we plan to include ED-RMF-1-pion production model.
-The STA-QMC calculation includes both single and two nucleon final states but at present it is only valid over a more restricted kinematic range and does not account for nuclear excitations.
+The STA-QMC calculation includes both single and two nucleon final states but is valid over a more restricted kinematic range and does not account for nuclear excitations.
 Extension of the theory to higher and lower values of momentum transfer are currently under way.
 In future communications we plan to report on the comparisons of response functions (including the nucleon resonance region) for $\rm{}^{12}C$, $\rm{}^{40}Ca$ and $\rm{}^{56}Fe$ to the predictions of ED-RMF, SuSAv2 and the genie MC generator.
-VIII Supplemental Materials
 The extracted ${\cal R}_{L}$ and ${\cal R}_{T}$ values (shown in the figures) are available as supplemental materials sup .
-In addition, the values of the Christy-Bodek fit to ${\cal R}_{L}$ and ${\cal R}_{T}$ over a larger range of $\nu$ (and also listing separately the contributions of nuclear excitations, quasielastic scattering, transverse enhancement, and pion production processes) are available as supplemental materials.
-VIII.1 Acknowledgments
+In addition, the values of the Christy-Bodek 2024 fit to ${\cal R}_{L}$ and ${\cal R}_{T}$ over a larger range of $\nu$ (and also listing separately the contributions of nuclear excitations, quasielastic scattering, transverse enhancement, and pion production processes) are available as supplemental materials.
+VIII Acknowledgments
 We thank Noah A.
@@
 Appendix A Comparison to SuSAv2 in bins of $\bf q$ and $Q^{2}$
+Comparisons to the predictions of SuSAv2 are shown in Figures 21-28.
+Appendix B Quasi-deuteron contribution
+The Quasi-deuteron contribution to the photoproduction ($Q^{2}=0$) cross section ($\sigma_{Quasi-D}$) as a function of $\nu=\varepsilon_{\gamma}$ is taken from Plujko et al.
+(2018).
+For a nucleus with $N$ neutrons, $Z$ protons and mass number $A=N+Z$, the $\sigma_{Quasi-D}$ is equal to
+$\sigma_{Quasi-D}(\varepsilon_{\gamma})=397.8\;\frac{NZ}{A}\frac{(\varepsilon_{\gamma}-2.224)^{3/2}}{\varepsilon_{\gamma}^{3}}\phi(\varepsilon_{\gamma})$ (42)
+with $\varepsilon_{\gamma}$ in MeV and $\sigma_{QD}$ in units of mb. The function $\phi(\varepsilon_{\gamma})$ accounts for the Pauli-blocking of the excited neutron-proton pair in the nuclear medium.
+$\phi(\varepsilon_{\gamma}<20\;{\rm MeV})={\rm exp}(-73.3/\varepsilon_{\gamma}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\$
+$\phi(20<\varepsilon_{\gamma}<140\;{\rm MeV})=8.3714\times 10^{-2}$
+$-9.8343\times 10^{-3}\varepsilon_{\gamma}+4.1222\times 10^{-4}\varepsilon_{\gamma}^{2}$
+$-3.4762\times 10^{-6}\varepsilon_{\gamma}^{3}+9.3537\times 10^{-9}\varepsilon_{\gamma}^{4},$
+$\phi(\varepsilon_{\gamma}>140\;{\rm MeV})={\rm exp}(-24.2/\varepsilon_{\gamma}).$ (43)
+Appendix C Supplemental Materials
+The following are available as supplemental materials sup
+(a) Comparison of the normalized $^{12}C$ electron scattering cross sections to the universal fit (15 Ffgures).
+(b) Tables of the extracted ${\cal R}_{L}$ and ${\cal R}_{T}$ values for bins of $\bf q$ and bins of $Q^{2}$.
+(c) Tables of the values of universal fit values of ${\cal R}_{L}$ and ${\cal R}_{T}$ (the contributions of nuclear excitations, quasielastic scattering, transverse enhancement, and pion production processes are listed separately).
 Figure 21: Comparison to ED-RMF and SuSAv2 for fixed bins in q: Our extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ for $\rm{}^{12}C$} for $\bf q$ values of 0.10, 0.148, 0.167, 0.205 and 0.240 GeV versus $\nu$.